### v2

# sample.bib

Abstract.
We generalize the classification of isomorphism classes of Schubert varieties in [RS21] in complete flag varieties $G/B$ to a class of partial flag varieties $G/P$. In particular, we classify all Schubert varieties in $G/P$ where $P$ is a minimal parabolic subgroup, and all Schubert surfaces which are two-dimensional Schubert varieties. We also obtain several classes of isomorphisms of Schubert varieties in flag varieties $G/P$ where $P$ is a maximal parabolic subgroup. In particular, we find an upper bound of the number of isomorphism classes of Schubert threefolds which are three-dimensional Schubert varieties.
1. Introduction
Schubert varieties form an extensively studied class of algebraic varieties whose properties are often characterized by combinatorics. The isomorphism problem for Schubert varieties, first raised by Develin, Martin, and Reiner in [DMR07], asks for a classification of all Schubert varieties up to algebraic isomorphism. In the same paper, they classified a class of smooth Schubert varieties in type $A$ partial flag varieties. Using Cartan equivalence, Richmond and Slofstra solved this problem for Schubert varieties in complete flag varieties in [RS21]. On the other hand, Richmond, Țarigradschi, and Xu solved this problem for cominuscule Schubert varieties in [RTX24] using labeled posets.
To describe our results, we set the following notations. We only consider Schubert varieties of finite types over $\mathbb{C}$ for simplicity. Let $G$ be a complex reductive Lie group, $T$ a maximal torus, and $B$ a Borel subgroup containing $T$. Then a root system is defined, with the correspondent set of simple reflections $S$, Cartan matrix $A=(a_{st})_{(s,t)\in S^{2}}$, and the Weyl group $W$. The pair $(W,S)$ forms a Coxeter system. A standard parabolic subgroup $P$ contains $B$ corresponding to a subset $I$ of $S$. We denote the flag variety $G/P$ by $X(A,I)$. The subset $I$ generates a subgroup $W_{I}\subset W$. In every coset $W/W_{I}$, we take the elements with minimal length and form a set $W^{I}$. Then a Schubert variety is the closure of the $B$ orbit $BwP/P$ for some $w\in W^{I}$, and we write this variety as $X(w,A,I)$ since it is uniquely determined by the triple $(w,A,I)$. We denote the support of $w$ to be
$S(w)=\{s\in S|s\leq w\}.$
In our notations, Richmond and Slofstra classified Schubert varieties of the form $X(w,A,\varnothing)$. Motivated by their methods, we extend their work.
Main Theorem 1.1 (See Proposition 4.5, Proposition 4.9, and Theorem 5.5).
Let
$A=(a_{st})_{(s,t)\in S^{2}}$ and $A^{\prime}=(a^{\prime}_{s^{\prime}t^{\prime}})_{(s^{\prime},t^{\prime})\in S^{\prime 2}}$
be two Cartan matrices with associated Weyl groups $W$ and $W^{\prime}$, and sets of simple reflections $S$ and $S^{\prime}$, respectively. Let $I\subset S$ and $I^{\prime}\subset S^{\prime}$. Take $w\in W^{I}$ and $w^{\prime}\in W^{\prime I^{\prime}}$. Assume that $|S(w)\cap I|\leq 1$ and $|S(w^{\prime})\cap I^{\prime}|\leq 1$. Then the following are equivalent:
  1. (1)
    the Schubert varieties $X(w,A,I)$ and $X(w^{\prime},A^{\prime},I^{\prime})$ are algebraically isomorphic;
  2. (2)
    there exists a bijection $\tau:S(w)\to S(w^{\prime})$ sending $S(w)\cap I$ to $S(w^{\prime})\cap I^{\prime}$, such that:
    1. (a)
      for some reduced word $w=s_{1}\cdots s_{k}$, $w^{\prime}=\tau(s_{1})\cdots\tau(s_{k})$ is also a reduced word;
    2. (b)
      for any $t_{1},t_{2}\in S(w)$, $a_{t_{1},t_{2}}=a^{\prime}_{\tau(t_{1}),\tau(t_{2})}$ whenever $t_{1}t_{2}\leq w$.
In particular, if $X(w,A,I)$ and $X(w^{\prime},A^{\prime},I^{\prime})$ are isomorphic, then $|S(w)\cap I|=|S(w^{\prime})\cap I^{\prime}|$.
Since we allow the subset $I$ to satisfy the condition $|S(w)\cap I|\leq 1$, a larger class of Schubert varieties is classified. It also completely solves the isomorphism problem of Schubert varieties in flag varieties corresponding to minimal parabolic subgroups, in which case, $|I|=1$.
Schubert varieties of dimension zero and one are isomorphic to a point and the projective line $\mathbb{P}^{1}$, respectively. Applying the main theorem, we can easily classify the Schubert varieties of dimension two without explicit computation.
The last statement of our main theorem gives a necessary condition for the isomorphism of Schubert varieties, which can be slightly generalized; see Proposition 4.5.
For two Schubert varieties $X(w,A,I)$ and $X(w^{\prime},A^{\prime},I^{\prime})$, if $|S(w)|=|S(w^{\prime})|$, then we say these two Schubert varieties are equally supported. The way we prove our criterion enables us to attain a sufficient condition (Theorem 5.5) to determine whether the given equally supported pair is an isomorphism.
It is also interesting to consider whether our sufficient condition is necessary. It is still unknown, and we will leave it for future study.
When the dimensions of Schubert varieties are greater than two, the isomorphic pairs that are not equally supported arise. We construct classes of such isomorphic pairs from folding the root systems; see Theorem 7.1. Essentially, the isomorphisms constructed in this way are isomorphisms of Schubert varieties in partial flag varieties corresponding to maximal parabolic subgroups; in other words, $|S\setminus I|=1$. Our theorem gives an upper bound of the number of isomorphism classes of three-dimensional Schubert varieties.
1.1. Outline of the Paper
In Section 2, we show some illustrative examples of Theorem 1.1. At the end of this section, we show the classification of Schubert surfaces; see Theorem 2.6. Then we discuss preliminaries in Section 3. The following two sections give a type-independent proof for our main theorem. In Section 4, we derive some necessary conditions of the isomorphism of Schubert varieties, and we prove that condition 1 implies condition 2 of the main theorem (Proposition 4.9). Our primary tool is the cohomology ring whose cup product can be computed by Chevalley’s formula in [FW01, Lemma 8.1]. We need more explicit computation and discussion than what was done in [RS21]. Next, in Section 5, we obtain a practical sufficient condition for equally supported Schubert varieties (Theorem 5.5), and, as a corollary, we prove that condition 2 implies condition 1 of the main theorem, finishing the proof of the main theorem. The next three sections focus on the non-equally supported pairs of Schubert varieties. In Section 6, we recall the notion of folding the root system by an automorphism. The literature does not fully state these materials, so we present them for complement. In Section 7, we state and prove Theorem 7.1, obtaining a criterion of isomorphisms of non-equally supported pairs of Schubert varieties. Finally, in Section 8, we specify Theorem 7.1 in different types. As an application, we compute an upper bound of the number of isomorphism classes of Schubert threefolds.
Acknowledgment
The author thanks Weihong Xu for introducing him to the isomorphism problem, providing him with relevant materials, and guiding him in writing the thesis. He also thanks Jing-Song Huang and Rui Xiong for their instruction on representation theory. He was partially supported by NSFC grant 12426507.
2. Classification of Schubert Surfaces
We provide several examples of Schubert varieties to illustrate our context. We will use $s_{i}$ to denote the simple reflection corresponding to the $i$-th row or column of the Cartan matrix.
Example 2.1.
Considering the three Schubert varieties
$X\left(s_{1}s_{2}s_{3}s_{4},\begin{pmatrix}2&-1&0&0\\ -1&2&-1&0\\ 0&-1&2&-1\\ 0&0&-1&2\end{pmatrix},\{s_{1}\}\right),$
$X\left(s_{1}s_{2}s_{3}s_{4},\begin{pmatrix}2&-1&0&0\\ -1&2&-1&0\\ 0&-1&2&-1\\ 0&0&-2&2\end{pmatrix},\{s_{1}\}\right),$
$X\left(s_{1}s_{2}s_{3}s_{4},\begin{pmatrix}2&-1&0&0\\ -1&2&-1&0\\ 0&-2&2&-1\\ 0&0&-1&2\end{pmatrix},\{s_{1}\}\right),$
of types $A_{4}$, $B_{4}$, and $F_{4}$, respectively. The upper triangular parts of three Cartan matrices are the same, so by the main theorem, they are isomorphic.
We call a two-dimensional Schubert variety a Schubert surface. For $n=1,2,3$, we let $M_{n}$ denote the Cartan matrix
$M_{n}=\begin{pmatrix}2&-1\\ -n&2\end{pmatrix}.$
In other words, $M_{1}$, $M_{2}$, and $M_{3}$ are the Cartan matrices of type $A_{2}$, $B_{2}$, and $G_{2}$, respectively. Note that $s_{1}$ and $s_{2}$ correspond to the long and the short simple root, respectively.
Example 2.2.
Consider the Schubert surfaces $X(s_{1}s_{2},M_{n},\{s_{1}\})$. Since the entries $a_{12}=-1$ of Cartan matrices are equal, by the main theorem, these three surfaces are isomorphic. Taking $n=1$, $X(s_{1}s_{2},M_{1},\{s_{1}\})$ is a Schubert surface in the flag variety $Gr(1,3)=\mathbb{P}^{2}$, which is also two dimensional. It follows that $X(s_{1}s_{2},M_{n},\{s_{1}\})$ is isomorphic to $\mathbb{P}^{2}$. Similarly, the Schubert surfaces $X(s_{1}s_{2},M_{n},\varnothing)$ are isomorphic. By [RS21, Example 1.1], they are isomorphic to the first Hirzebruch surface $\Sigma_{1}$.
Example 2.3.
Consider the Schubert surfaces $X(s_{2}s_{1},M_{n},\{s_{2}\})$. The corresponding entries $a_{21}=-n$, so these three varieties are not isomorphic to each other. Note that geometrically, $X(s_{2}s_{1},M_{1},\{s_{2}\})$ is the same as $X(s_{1}s_{2},M_{1},\{s_{1}\})$, so it is isomorphic to $\mathbb{P}^{2}$. On the other hand, it can be shown that $X(s_{2}s_{1},M_{2},\{s_{2}\})$ is the cone over a smooth conic. In fact, the flag variety $X(s_{2}s_{1},M_{2})$ is a smooth quadric hypersurface in $\mathbb{P}^{4}$ defined by the equation $x_{1}x_{5}+x_{2}x_{4}+x_{3}^{2}=0$, and the Schubert variety $X(s_{2}s_{1},M_{2},\{s_{2}\})$ lies in the hyperplane $x_{5}=0$. Hence the Schubert surface is defined by the equation $x_{2}x_{4}+x_{3}^{2}=0$ in $\mathbb{P}^{3}$, which is the cone over a smooth conic.
Example 2.4.
Consider the Schubert surfaces $X(s_{2}s_{1},M_{n},\varnothing)$. The corresponding entries $a_{21}=-n$, so these four surfaces are not isomorphic to each other. As mentioned in [RS21, Example 1.1], $X(s_{2}s_{1},M_{n},\varnothing)$ is isomorphic to the $n$-th Hirzebruch surface $\Sigma_{n}$.
Example 2.5.
Consider the Schubert surface $X(s_{1}s_{2},2I_{2},\varnothing)$, where $I_{2}$ is the $2\times 2$ identity matrix. The corresponding entry $a_{12}=0$, so it differs from any Schubert surfaces mentioned above. The Cartan matrix is of type $A_{1}\times A_{1}$, so the flag variety can be identified with $\mathbb{P}^{1}\times\mathbb{P}^{1}$. Comparing the dimensions, we conclude that $X(s_{1}s_{2},2I_{2},\varnothing)$ is $\mathbb{P}^{1}\times\mathbb{P}^{1}$.
The Schubert surfaces in the above four examples are all possible Schubert surfaces. To be more precise, we have the following theorem.
Theorem 2.6 (Classification of Schubert Surfaces).
There are seven isomorphism classes of Schubert surfaces. Precisely, any Schubert surface is isomorphic to one and exactly one of the following varieties:
  1. (1)
    product of projective lines $\mathbb{P}^{1}\times\mathbb{P}^{1}$,
  2. (2)
    projective plane $\mathbb{P}^{2}$,
  3. (3)
    the $n$-th Hirzebruch surface $\Sigma_{n}$ for $n=1,2,3$,
  4. (4)
    the cone over a smooth conic, and
  5. (5)
    the variety $X(s_{2}s_{1},M_{3},\{s_{2}\})$.
Proof.
Recall that the dimension of a Schubert variety $X(w,A,I)$ is the length $\ell(w)$ of $w\in W^{I}$ so that we can write $w=st\in W^{I}$ as the product of two simple reflections $s$ and $t$. Lemma 3.9 (or [RS16, Lemma 4.8]) implies that, without loss of generality, we can assume that $S(w)=S$. In other words, $S=\{s,t\}$ and $A$ has rank two. Since a root system of rank two have type $A_{1}\times A_{1}$, $A_{2}$, $B_{2}$, or $G_{2}$, after appropriately indexing rows and columns, $A=2I_{2}$ or $A=M_{n}$ for $n=1,2,3$. $I=\{s\}$ or $\varnothing$. Then there are only finitely many possibilities for $X(w,A,I)$, which we write in the following classes:
  1. (1)
    $X(s_{1}s_{2},2I_{2},\varnothing)$, isomorphic to $\mathbb{P}^{1}\times\mathbb{P}^{1}$;
  2. (2)
    $X(s_{1}s_{2},M_{n},\varnothing)(n=1,2,3)$ and $X(s_{2}s_{1},M_{1},\varnothing)$, isomorphic to $\Sigma_{1}$;
  3. (3)
    $X(s_{2}s_{1},M_{n},\varnothing)(n=2,3)$, isomorphic to $\Sigma_{n}$;
  4. (4)
    $X(s_{1}s_{2},M_{n},\{s_{1}\})(n=1,2,3)$ and $X(s_{2}s_{1},M_{1},\{s_{2}\})$, isomorphic to $\mathbb{P}^{2}$;
  5. (5)
    $X(s_{2}s_{1},M_{2},\{s_{2}\})$, isomorphic to the cone over a smooth conic;
  6. (6)
    $X(s_{2}s_{1},M_{3},\{s_{2}\})$.
By the last statement of Theorem 1.1, any variety in the first three classes is not isomorphic to a variety in the remaining three classes. In classes 1, 2, and 3, the corresponding entry of the Cartan matrix is 0, $-1$, and $-n(n=2,3)$, so there are four isomorphic classes. In classes 4, 5, and 6, the entries of the Cartan matrices are $-1$, $-2$, and $-3$, so there are three more isomorphic classes. Therefore, the proof is complete. ∎
3. Preliminaries
3.1. The Weyl Group and the Root System
Given a complex reductive group $G$ with a maximal torus $T$ and a Borel subgroup $B$ containing $T$, a root system $R$ is determined. Let $R=R^{+}\cup R^{-}$ be the choice of positive/negative roots determined by $B$, and $\Delta$ the set of simple roots. With a choice of ordered for $\Delta$, the root system $R$ defines a Cartan matrix
$A=(a_{st})_{(s,t)\in S^{2}},$ where $\mbox{where }a_{st}=\langle\alpha_{s}^{\vee},\alpha_{t}\rangle=\frac{2\langle\alpha_{s},\alpha_{t}\rangle}{\langle\alpha_{s},\alpha_{s}\rangle}.$
Any positive root $\alpha$ defines a reflection $s_{\alpha}$; conversely, a reflection $s$ defines a positive root $\alpha_{s}$. The set of simple reflections $S$ generates the Weyl group $W$, which can be identified with $W=N(T)/T$. The pair $(W,S)$ is a Coxeter system. For any Weyl group element $w\in W$, we denote
$S(w)=\{s\in S|s\leq w\}$
to be the support of $w$. If $S(w)=S$, then we say that $w$ is fully supported. The element $w$ can be written as a product of simple reflections, and if $w=s_{1}\cdots s_{k}$ with $k$ minimal, then we say the sequence $(s_{1},\cdots,s_{k})$ (or the product $s_{1}\cdots s_{k}$) is a reduced word and denote the set of reduced words to be $RW(w)$. The Coxeter length function $\ell:W\to\mathbb{Z}_{\geq 0}$ is defined to be the length of any reduced word. There is a unique longest element, denoted by $w_{0}$. We define the right descent set of $w\in W$ to be
$D_{R}(w)=\{s\in S|\ell(us)=\ell(u)-1\}.$
In other words, $D_{R}(w)$ is the set of simple reflections $s$ that are the rightmost elements in some reduced words of $w$.
A standard parabolic subgroup $P$ corresponds to a subset $I\subset S$. We let the subgroup $W_{I}\subset W$ be generated by $I$, and the set $W^{I}$ to be the representatives of $W/W_{I}$ with minimal length. It is easy to see that a simple reflection $s\in D_{R}(w)$ if and only if $w\notin W^{\{s\}}$. For $w\in W$, We let $w^{\min}\in W^{I}$ be the minimal length representative in the coset $wW_{I}$. On the other hand, $I$ defines a subset of simple roots $\Delta_{I}$. We write the root subsystem spanned by $\Delta_{I}$ as $R_{I}$.
On the Weyl group $W$ (or, more generally, a Coxeter group), there is a partial order called the Bruhat order, defined by the following equivalent conditions. The equivalence is a restatement of [Hum92, Theorem 5.10].
Lemma 3.1.
Suppose that $u,v\in W$. Then the following are equivalent:
  1. (a)
    For any reduced word $v=s_{1}\cdots s_{k}$ for $v$, $u=s_{i_{1}}\cdots s_{i_{l}}$;
  2. (b)
    There exists a sequence $u_{0}=u,u_{1},\cdots,u_{n}=v$, such that $u_{i}=u_{i-1}t$ for some reflection $t$ and $\ell(u_{i})>\ell(u_{i-1})$.
For simplicity, we write $[u,v]^{I}=[u,v]\cap W^{I}$. [BB05, Proposition 2.5.1] shows the key property of Bruhat order on $W^{I}$. We state as follows.
Lemma 3.2.
The map $w\mapsto w^{\min}$ is order-preserving.
Now we consider the interaction between Weyl groups and root systems. We define the inversion set of $w\in W$ to be
$I(w)=R^{+}\cap w(R^{-}).$
The following statement is part of [Hum92, Lemma 1.6].
Lemma 3.3.
Let $w\in W$ and $s\in S$. Then $\ell(sw)>\ell(w)$ if and only if $w^{-1}(\alpha_{s})\in R^{+}$. Similarly, we have $\ell(ws)>\ell(w)$ if and only if $w(\alpha_{s})\in R^{+}$.
As a corollary, if $w\in W^{I}$, then $\ell(ws)>\ell(w)$ for all $s\in I$, which implies $w(\alpha_{s})\in R^{+}$. It follows that $w(R_{I}^{+})\subset R^{+}$, or $R_{I}^{+}\subset R^{+}\cap w^{-1}(R^{+})$. In other words, we have the following lemma.
Lemma 3.4.
Let $w\in W^{I}$. Then $I(w^{-1})\subset R^{+}\setminus R_{I}^{+}$.
3.2. Lie Algebras
Denote the Lie algebra of $G$ and $T$ as $\mathfrak{g}$ and its Cartan subalgebra $\mathfrak{h}$, respectively. Each coroot $\alpha^{\vee}=\frac{2\alpha}{\langle\alpha,\alpha\rangle}$ can be identify as an element $h_{\alpha}$ in $\mathfrak{h}$, and $\{h_{\alpha}\}_{\alpha\in\Delta}$ is a basis of $\mathfrak{h}$. We denote the root space corresponding to $\alpha$ as $\mathfrak{g}_{\alpha}$. Then we have the root space decomposition
$\mathfrak{g}=\mathfrak{h}\oplus\bigoplus_{\alpha\in R}\mathfrak{g}_{\alpha}.$
We also denote the nilpotent Lie algebras
$\mathfrak{n}^{\pm}=\bigoplus_{\alpha\in R^{\pm}}\mathfrak{g}_{\alpha},$
and the Borel subalgebra
$\mathfrak{b}=\mathfrak{h}\oplus\mathfrak{n}^{+}.$
Write $e_{\alpha}$ and $f_{\alpha}$ to be nonzero elements in $\mathfrak{g}_{\alpha}$ and $\mathfrak{g}_{-\alpha}$ for any positive root $\alpha$, respectively. Then $\{e_{\alpha}\}_{\alpha\in\Delta}$ and $\{f_{\alpha}\}_{\alpha\in\Delta}$ generate $\mathfrak{n}^{+}$ and $\mathfrak{n}^{-}$, respectively. For $w\in W$, we denote
$\mathfrak{n}^{+}_{w}=\bigoplus_{\alpha\in I(w)}\mathfrak{g}_{\alpha},$
which is a nilpotent Lie algebra.
Let $\lambda$ be a dominant integral weight. Then there is a unique irreducible $\mathfrak{g}$-module $L_{\lambda}$ with the highest weight $\lambda$. It can be constructed in the following way. Let $\mathbb{C}_{\lambda}$ be a one-dimensional complex vector space spanned by $\omega=\omega_{\lambda}$, whose $\mathfrak{b}$-module structure is given by
$h\omega$ $=\lambda(h)\omega,$ $h\in\mathfrak{h},$
$e\omega$ $=0,$ $e\in\mathfrak{n}^{+}.$
Then the Verma module with highest weight $\lambda$ is
$M_{\lambda}=U(\mathfrak{g})\otimes_{U(\mathfrak{b})}\mathbb{C}_{\lambda},$
where $U(\mathfrak{g})$ and $U(\mathfrak{b})$ are the universal enveloping algebras of $\mathfrak{g}$ and $\mathfrak{b}$, respectively. It has a submodule, denoted by $M^{1}_{\lambda}$, which is generated by $f_{\alpha}^{\lambda(h_{\alpha})+1}\otimes 1=f_{\alpha}^{\lambda(h_{\alpha})+1}\omega$. Then by [Kir08, Theorem 8.28],
$L_{\lambda}=M_{\lambda}/M^{1}_{\lambda}.$
3.3. Schubert Varieties
The variety $G/P$, called the (partial) flag variety, is determined by the Cartan matrix $A$ and the subset $I$, so we denote it as $X(A,I)$.
We say a weight $\lambda$ is an $I$-regular weight if it is a dominant integral weight and satisfies, for a simple reflection $s\in S$, $\lambda(\alpha_{s}^{\vee})=0$ if and only $s\in I$. The following result is in [Kum12, Lemma 7.1.2].
Lemma 3.5.
Suppose that $\lambda$ is an $I$-regular weight, and $V$ is a finite-dimensional highest weight module with the highest weight vector $\omega$. Then there is a closed immersion
$G/P\hookrightarrow\mathbb{P}(V),$ $gP\mapsto[g\omega],$
where $[g\omega]$ denote the line through $g\omega$.
A Schubert cell in $G/P$ is a subvariety $BwP/P$. To be more precise, we need to replace $w$ with a representative in $N(T)$, but any such choice gives the same variety. Different choices of $w$ may give the same Schubert variety. In fact, $BwP/P$ only depends on the coset $wW_{I}$, so we usually choose the minimal representative. If we take $w\in W^{I}$, then as varieties, $BwP/P\cong\mathbb{A}^{\ell(w)}$.
A Schubert variety in $G/P$, denoted by $X(w,A,I)$, is the Zariski closure of the Schubert cell $BwP/P$. The triple $(w,A,I)$ determines the Schubert variety. $X(w,A,I)$ has an affine stratification. If we take $w\in W^{I}$, then we have a disjoint union
$X(w,A,I)=\bigcup_{u\in[1,w]^{I}}BuP/P.$
Definition 3.6.
There is a $\mathbb{Z}$-basis $\sigma_{v}$, indexed by $v\in[1,w]^{I}$, for the integral cohomology group $H^{*}(X(w,A,I))$. We call this basis to be Schubert basis, denoted by $\Sigma(w,A,I)$. The collection of Schubert basis elements of degree two is denoted by $\Sigma_{1}(w,A,I)$.
The following lemma is essentially [RS21, Proposition 4.1], where the case $I=\varnothing$ was proven. Nevertheless, the statement holds in a more general case, and the proof can be extended to our context when replacing $X(w,A)$ and $X(w^{\prime},A^{\prime})$ in their paper by $X(w,A,I)$ and $X(w^{\prime},A^{\prime},I^{\prime})$, respectively.
Lemma 3.7.
Suppose that $f:X(w,A,I)\to X(w^{\prime},A^{\prime},I^{\prime})$ is an isomorphism of algebraic varieties. Then the induced map
$\varphi^{*}:H^{*}(X(w,A,I))\to H^{*}(X(w^{\prime},A^{\prime},I^{\prime}))$
is a graded ring isomorphism that identifies Schubert bases $\Sigma(w,A,I)\to\Sigma(w^{\prime},A^{\prime},I^{\prime})$ and $\Sigma_{1}(w,A,I)\to\Sigma_{1}(w^{\prime},A^{\prime},I^{\prime})$.
In other words, both $\Sigma(w,A,I)$ and $\Sigma_{1}(w,A,I)$ are determined by the variety structure of the Schubert variety $X(w,A,I)$. From this definition, we have an identification
$i:[1,w]^{I}\to\Sigma(w,A,I),$ $u\mapsto\sigma_{u}.$
Each basis element $\sigma_{u}$ has degree $2\ell(u)$, written by $\deg\sigma_{u}=2\ell(u)$, i.e.,
$\sigma_{u}\in H^{2\ell(u)}(X(w,A,I)).$
Recall that the set of simple reflections in $[1,w]^{I}$ is just $S(w)\setminus I$. Hence the restriction of $i$ gives a one-one correspondence
$i:S(w)\setminus I\to\Sigma_{1}(w,A,I),$ $s\mapsto\sigma_{s}.$
The cup product of $H^{*}(X(w,A,I))$ is given by the Chevalley’s formula in [FW01, Lemma 8.1].
Lemma 3.8 (Chevalley’s formula).
Suppose that $\alpha\in\Delta\setminus\Delta_{I}$ and $w\in W^{I}$. Then
$\sigma_{s_{\alpha}}\sigma_{w}=\sum\omega_{\alpha}(\beta^{\vee})\sigma_{(ws_{\beta})^{\min}},$
the sum over all positive roots $\beta$ such that $\ell((ws_{\beta})^{\min})=\ell(w)+1$. Here, $\omega_{\alpha}$ is the fundamental weight corresponding to $\alpha$.
Remark.
If $\ell((ws_{\beta})^{\min})=\ell(w)+1$, then either $ws_{\beta}\leq w$ or $ws_{\beta}>w$. If the former one holds, then we have $(ws_{\beta})^{\min}\leq w^{\min}=w$, which is impossible. Hence $ws_{\beta}>w$.
When considering the variety structure of a Schubert variety, it is usually convenient to embed it into a smaller flag variety. This is given by [RS16, Lemma 4.8].
Lemma 3.9.
We write $A_{w}=(a_{st})_{(s,t)\in S(w)^{2}}$ to be the Cartan submatrix of $A$. Then the inclusion $X(A_{w},S(w)\cap I)\hookrightarrow X(A,I)$ restricts to an isomorphism
$X(w,A_{w},S(w)\cap I)\cong X(w,A,I).$
Remark.
The Weyl group determined by $A_{w}$ is $W_{S(w)}$, and $w\in(W_{S(w)})^{S(w)\cap I}$, so the Schubert variety $X(w,A_{w},S(w)\cap I)$ is well-defined.
4. The Cohomology Rings of Schubert Varieties
In this section, we will use the cohomology ring to study the intrinsic properties of a Schubert variety. In other words, we will obtain combinatorial information from an isomorphism class of the Schubert variety. Some results have been proven in [RS21, Section 4] for the case $I$ being an empty set, and the proofs are generalized. In the following, we fix a Schubert variety $X(w,A,I)$ with $w\in W^{I}$. Due to Lemma 3.9, we assume that $S=S(w)$.
4.1. General Situation
First, we can reconstruct the poset $[1,w]^{I}$. For any
$\sigma=\sum_{v\in W^{I}}a_{v}\sigma_{v}\in H^{*}(X(w,A,I)),$
we define its support, denoted by $\Sigma(\sigma)$, to be the collection of $\sigma_{v}$ with $a_{v}\neq 0$. Let $\prec$ be the partial order on $\Sigma(w,A,I)$, the Schubert basis, generated by the relation $\sigma_{u}\prec\sigma_{v}$ if $\sigma_{v}\in S(\sigma_{s}\sigma_{u})$ for some $\sigma_{s}\in\Sigma_{1}(w,A,I)$. By Lemma 3.7, the poset $\Sigma(w,A,I)$ is determined by the variety structure of $X(w,A,I)$.
Lemma 4.1.
The identification $i:[1,w]^{I}\to\Sigma(w,A,I)$ is a poset isomorphism.
Proof.
It is sufficient to prove that, for $u,v\in[1,w]^{I}$, $u<v$ if and only if $\sigma_{u}\prec\sigma_{v}$, and it suffices to consider the covering relations. If $u<v$ and $\ell(v)=\ell(u)+1$, then $v=us_{\beta}$ for some positive root $\beta$. Since $u,v\in W^{I}$, $\beta\notin R_{I}$. Hence there exists a simple root $\alpha\in\Delta\setminus\Delta_{I}$ s.t. $\omega_{\alpha}(\beta^{\vee})\neq 0$. It follows that $\sigma_{v}\in\Sigma(\sigma_{s_{\alpha}}\sigma_{u})$.
Conversely, if $\sigma_{v}\in\Sigma(\sigma_{s}\sigma_{u})$ for some $\sigma_{s}\in\Sigma_{1}(w,A,I)$, then $\sigma_{v}=\sigma_{(us_{\beta})^{\min}}$ for some positive root $\beta$. By the remark below Lemma 3.9, we have $us_{\beta}>u$, and it follows that $u<v$. ∎
Given a subset $J^{\prime}$ of $\Sigma_{1}(w,A,I)$, let $H^{J^{\prime}}$ be the subring of $H^{*}(X(w,A,I))$ generated by
$\sigma_{s}\in\Sigma_{1}(w,A,I)\setminus J^{\prime},$
and let
$\Sigma(w,A,I)^{J^{\prime}}=\bigcup_{\sigma\in H^{J^{\prime}}}\Sigma(\sigma).$
Lemma 4.2.
Suppose that there is a subset $J\subset S\setminus I$. Then the poset isomorphism $i:[1,w]^{I}\to\Sigma(w,A,I)$ restricts to another poset isomorphism
$i^{J}:[1,w]^{I\cup J}\to\Sigma(w,A,I)^{i(J)}.$
Proof.
It suffices to prove that $i^{J}$ is a bijection. First, we show that im $\mbox{im }i^{J}\subset\Sigma(w,A,I)^{i(J)}$ by induction on length. Suppose that $v\in[1,w]^{I\cup J}$ and $v=s_{1}\cdots s_{k}$ is a reduced word. Let $u=s_{1}v=s_{2}\cdots s_{k}\in[1,w]^{I\cup J}$. Then since $\ell(u)<\ell(v)$, by induction, $\sigma_{u}\in\Sigma(w,A,I)^{i(J)}$. Apply Lemma 3.3, we conclude that $u^{-1}(\alpha_{s_{1}})\in R^{+}$. Then
$u^{-1}(\alpha_{s_{1}})=v^{-1}s_{1}(\alpha_{s_{1}})=v^{-1}(-\alpha_{s_{1}})\in I(v^{-1}).$
Since $v\in W^{I\cup J}$, by Lemma 3.4 we have $I(v^{-1})\subset R^{+}\setminus R_{I\cup J}^{+}$. Thus, there exists a simple root $\alpha\in\Delta\setminus\Delta_{I\cup J}$, such that $\omega_{\alpha}(u^{-1}(\alpha_{s_{1}}^{\vee}))\neq 0$. Then
$\sigma_{v}\in\Sigma(\sigma_{s_{\alpha}}\sigma_{u})\subset\Sigma(w,A,I)^{i(J)}$
since $u^{-1}v=u^{-1}s_{1}u$ is the reflection corresponding to the positive root $u^{-1}(\alpha_{s_{1}})$. Hence the inclusion im $\mbox{im }i^{J}\subset\Sigma(w,A,I)^{i(J)}$ holds.
Conversely, suppose that $u\in[1,w]^{I\cup J}$, $\alpha\in\Delta\setminus\Delta_{I\cup J}$ and $\sigma_{v}\in\Sigma(\sigma_{s_{\alpha}}\sigma_{u})$. We need to show that $v\in W^{I\cup J}$. Then the lemma follows from the similar induction process. If $v\notin W^{I\cup J}$, then there exists a reduced word $v=s_{1}\cdots s_{k}$ with $s_{k}\in{I\cup J}$. Since $u<v$ and $u\in W^{I\cup J}$, $u$ must have the form $u=s_{1}\cdots s_{k-1}$. But $\omega_{\alpha}(\alpha_{s_{k}}^{\vee})=0$, which contradicts to the assumption $\sigma_{v}\in\Sigma(\sigma_{s_{\alpha}}\sigma_{u})$. ∎
Now we show how to reconstruct the right descent set. We define the right descent set of $\sigma_{v}$ to be
$D_{R}(\sigma_{v})=\left\{\sigma_{s}\in\Sigma_{1}(w,A,I)|\sigma_{v}\notin\Sigma(w,A,I)^{\{\sigma_{s}\}}\right\}.$
Lemma 4.3.
Suppose that $v\in[1,w]^{I}$. Then the poset isomorphism $i:[1,w]^{I}\to\Sigma(w,A,I)$ restricts to a bijection $i_{v}:D_{R}(v)\to D_{R}(\sigma_{v})$.
Proof.
Recall that $s\in D_{R}(v)$ if and only if $v\notin W^{\{s\}}$. Since $v\in W^{I}$, $s\in D_{R}(v)$ if and only if $v\notin W^{I\cup\{s\}}$. Now the result follows from Lemma 4.2. ∎
The following lemma will be used to reconstruct the reduced words.
Lemma 4.4.
Suppose that $\sigma_{v}\in\Sigma(w,A,I)$ and $\sigma_{t}\in D_{R}(\sigma_{v})$. Then there exists a unique maximal element $\sigma_{u}\prec\sigma_{v}$ such that $\sigma_{u}\in\Sigma(w,A,I)^{\{\sigma_{t}\}}$.
Proof.
Under the identification $i$, the element $u$ satisfies $u<v$ and $u\in W^{I\cup\{t\}}$. Then by Lemma 3.2, $u\leq u_{0}$, the minimal representative of $v$ in the coset of $W/W_{I\cup\{t\}}$. But $u_{0}$ also satisfies $u_{0}<v$ and $u_{0}\in W^{I\cup\{t\}}$. Hence the unique maximal $\sigma_{u}$ is $\sigma_{u_{0}}$. ∎
We obtain the following proposition as a corollary, which gives a necessary condition.
Proposition 4.5.
If $S(w)\cap I$ is empty but $S(w^{\prime})\cap I^{\prime}$ is not, then $X(w,A,I)$ and $X(w^{\prime},A^{\prime},I^{\prime})$ are not isomorphic as algebraic varieties.
Proof.
Using Lemma 3.9, we can assume that $w,w^{\prime}$ are fully supported. Hence $I=\varnothing\neq I^{\prime}$. We can find two simple reflections $s^{\prime},t^{\prime}\in S^{\prime}$ such that $s^{\prime}t^{\prime}\in(1,w^{\prime}]^{I^{\prime}}$ and $s^{\prime}\in I$. In fact, taking a reduced word $w^{\prime}=s^{\prime}_{1}\cdots s^{\prime}_{k}$, we let $s^{\prime}=s^{\prime}_{i}$, where $i$ is largest integer such that $s^{\prime}_{i}\in I^{\prime}$, and $t^{\prime}=s^{\prime}_{j}$ with $j>i$ such that $s^{\prime}_{i}$ and $s^{\prime}_{j}$ cannot commute. Then $s^{\prime}t^{\prime}=s^{\prime}_{i}s^{\prime}_{j}\in(1,w^{\prime}]^{I^{\prime}}$. The integer $j$ exists since otherwise, $s_{i}$ commutes with all $s^{\prime}_{j}$ for all $j>i$, which implies
$w^{\prime}=s^{\prime}_{1}\cdots\widehat{s^{\prime}_{i}}\cdots s^{\prime}_{k}s^{\prime}_{i}\notin W^{\prime I^{\prime}},$
contradicting to the assumption that $w^{\prime}\in W^{\prime I^{\prime}}$.
Suppose that $X(w,A,\varnothing)$ and $X(w^{\prime},A^{\prime},I^{\prime})$ are algebraically isomorphic. Then the set $\Sigma_{1}(w,A,I)$ (resp. the poset $\Sigma(w,A,I)$) can be identified with the set $\Sigma_{1}(w^{\prime},A^{\prime},I^{\prime})$ (resp. the poset $\Sigma(w^{\prime},A^{\prime},I^{\prime})$). Hence there is a pair
$(\sigma_{t},\sigma_{st})\in\Sigma_{1}(w,A,I)\times\Sigma(w,A,I)$
of Schubert basis elements corresponding to the pair
$(\sigma_{t^{\prime}},\sigma_{s^{\prime}t^{\prime}})\in\Sigma_{1}(w^{\prime},A^{\prime},I^{\prime})\times\Sigma(w^{\prime},A^{\prime},I^{\prime}).$
Consider the unique maximal element $\sigma_{u}\prec\sigma_{st}$ (resp. $\sigma_{u^{\prime}}\prec\sigma_{s^{\prime}t^{\prime}}$) such that $\sigma_{u}\in\Sigma(w,A,I)^{\{\sigma_{t}\}}$ (resp. $\sigma_{u^{\prime}}\in\Sigma(w^{\prime},A^{\prime},I^{\prime})^{\{\sigma_{t^{\prime}}\}}$), which exists by Lemma 4.4. But $u=s$ and $u^{\prime}=1$, so
$\deg\sigma_{u}=2\neq 0=\deg\sigma_{u^{\prime}},$
which is a contradiction. ∎
Now we compute some entries of the Cartan matrix from the cohomology ring.
Lemma 4.6.
Let $s\in S$ and $t\in S\setminus I$ with $s\neq t$. If $st\in[t,w]^{I}$, then the coefficient of $\sigma_{st}$ in $\sigma_{t}^{2}$ is $-a_{st}$.
Proof.
By Chevalley’s formula, the coefficient of $\sigma_{st}$ in $\sigma_{t}^{2}$ is
$\omega_{\alpha_{t}}(t(\alpha_{s}^{\vee}))=\omega_{\alpha_{t}}(\alpha_{s}^{\vee}-a_{st}\alpha_{t}^{\vee})=-a_{st}.$
Lemma 4.7.
Let $r,s\in S$ and $t\in S\setminus I$ with $rst\in[1,w]^{I}$ being a reduced word. Then the coefficient of $\sigma_{rst}$ in $\sigma_{t}\sigma_{st}$ is $\delta_{rt}-a_{rt}+a_{rs}a_{st}$, where $\delta_{rt}=1$ if $r=t$, and $\delta_{rt}=0$ if $r\neq t$.
Proof.
By Chevalley’s formula, the coefficient of $\sigma_{rst}$ in $\sigma_{t}\sigma_{st}$ is
$\omega_{\alpha_{t}}(ts(\alpha_{r}^{\vee}))$ $=\omega_{\alpha_{t}}(t(\alpha_{r}^{\vee}-a_{rs}\alpha_{s}^{\vee}))=\omega_{\alpha_{t}}(\alpha_{r}^{\vee}-a_{rs}\alpha_{s}^{\vee}-(a_{rt}-a_{rs}a_{st})\alpha_{t}^{\vee})$
$=\delta_{rt}-(a_{rt}-a_{rs}a_{st})=\delta_{rt}-a_{rt}+a_{rs}a_{st}.$
4.2. Case $I=\{s\}$
We further assume that $I=\{s\}$. We are primarily interested in this situation.
First, we define the reduced word of $\sigma_{v}\in\Sigma(w,A,I)$. We let $\sigma_{s}$ be a symbol and define
$\widetilde{\Sigma}_{1}(w,A,I)=\Sigma_{1}(w,A,I)\cup\{\sigma_{s}\},$
and extend the bijection $S\setminus I\to\Sigma_{1}(w,A,I)$ to a bijection $S\to\widetilde{\Sigma}_{1}(w,A,I)$ by sending $s$ to the symbol $\sigma_{s}$. We define the reduced words of $\sigma_{v}\in\Sigma(w,A,I)$ inductively. First, define the reduced word of $\sigma_{1}\in H^{0}(X(w,A,\{s\}))$ to be the singleton of the empty word. Next, suppose that $\sigma_{v}\in\Sigma(w,A,I)$ with $\sigma_{t}\in D_{R}(\sigma_{v})$. Take $\sigma_{u}$ to be the unique maximal element described in Lemma 4.4, and assume that $\deg\sigma_{v}-\deg\sigma_{u}=2n$ for $n\in\mathbb{N}$. Then we define the sequence
$(\sigma_{s_{1}},\cdots,\sigma_{s_{k}},\cdots,\sigma_{t},\sigma_{s},\sigma_{t},\sigma_{s})$
is a reduced of $\sigma_{v}$, where $(\sigma_{s_{1}},\cdots,\sigma_{s_{k}})$ is a reduced word of $\sigma_{u}$, and the remaining part $(\cdots,\sigma_{t},\sigma_{s},\sigma_{t},\sigma_{s})$ of the sequence is an alternating sequence of $\sigma_{t}$ and $\sigma_{s}$ of length $n$. We also denote $RW(\sigma_{v})$ as the set of reduced words of $\sigma_{v}$, which is nonempty. The following lemma is clear.
Lemma 4.8.
Suppose that $\sigma_{v}\in\Sigma(w,A,I)$. Then the bijection $S\to\widetilde{\Sigma}_{1}$ induces an inclusion
$RW(\sigma_{u})\hookrightarrow RW(u),$ $(\sigma_{s_{1}},\cdots,\sigma_{s_{k}})\mapsto(s_{1},\cdots,s_{k}).$
The above lemmas allow us to prove the following proposition.
Proposition 4.9.
Let $A=(a_{st})_{(s,t)\in S^{2}}$ and $A^{\prime}=(a^{\prime}_{s^{\prime}t^{\prime}})_{(s^{\prime},t^{\prime})\in S^{\prime 2}}$ be two Cartan matrices with associated Weyl groups $W$ and $W^{\prime}$, and sets of simple reflections $S$ and $S^{\prime}$, respectively. Let $s\in S$ and $s^{\prime}\in S^{\prime}$. Take two fully supported elements $w\in W^{\{s\}}$ and $w^{\prime}\in W^{\prime\{s^{\prime}\}}$. If $X(w,A,\{s\})$ and $X(w^{\prime},A^{\prime},\{s^{\prime}\})$ are algebraically isomorphic, then there is a bijection $\tau:S\to S^{\prime}$ sending $s$ to $s^{\prime}$ such that:
  1. (a)
    for some reduced word $w=s_{1}\cdots s_{k}$, $w^{\prime}=\tau(s_{1})\cdots\tau(s_{k})$ is also a reduced word;
  2. (b)
    for any $t_{1},t_{2}\in S$, $a_{t_{1}t_{2}}=a^{\prime}_{\tau(t_{1})\tau(t_{2})}$ whenever $t_{1}t_{2}\leq w$.
Proof.
To simplify notations, we write $I=\{s\}$ and $I^{\prime}=\{s^{\prime}\}$. We denote $i^{\prime}$ as the identification $[1,w^{\prime}]^{I^{\prime}}\to\Sigma(w^{\prime},A^{\prime},I^{\prime})$. The bijection $\tau$ is taken to be the composition of bijections
$S\stackrel{{\scriptstyle i}}{{\longrightarrow}}\Sigma_{1}(w,A,I)\longrightarrow\Sigma_{1}(w^{\prime},A^{\prime},I^{\prime})\stackrel{{\scriptstyle i^{\prime-1}}}{{\longrightarrow}}S^{\prime}.$
Then $\tau$ identifies the reduced words by Lemma 4.8. Hence (a) is proven.
Now we check that (b) holds valid.
Suppose that $t_{1},t_{2}\in S$ such that $t_{1}t_{2}\leq w$. If $t_{1}=t_{2}$, then $a_{t_{1}t_{2}}=2=a^{\prime}_{\tau(t_{1})\tau(t_{2})}$. If $t_{1}t_{2}\in(1,w]^{I}$, then Lemma 4.6 implies that the coefficient of $\sigma_{t_{1}t_{2}}$ and $\sigma_{\tau(t_{1})\tau(t_{2})}$ in $\sigma_{t_{2}}^{2}$ and $\sigma_{\tau(t_{2})}^{2}$ are $-a_{t_{1}t_{2}}$ and $-a^{\prime}_{\tau(t_{1})\tau(t_{2})}$, respectively. Hence $a_{t_{1}t_{2}}=a^{\prime}_{\tau(t_{1})\tau(t_{2})}$.
Suppose that $t_{1}t_{2}\notin[1,w]^{I}$ for the remaining case. Then $t_{1}t_{2}=rs$ for some $r\in S\setminus I$. Taking a reduced word $w=s_{1}\cdots s_{k}$ corresponding to a reduced word of $\sigma_{w}$, let $i$ be the maximal integer such that $s_{i}=s$. We further assume that the reduced word of $w$ is taken to satisfy that $i$ is maximal among all such reduced words. Then $t=s_{i+1}\in S\setminus I$ satisfies $st\neq ts$. If $rst$ is not reduced, then $rst=s$ since $rst\neq r$ or $t$. It follows that $r=t$ and hence $m_{st}=2$, which contradicts the assumption that $s$ and $t$ cannot commute. Thus, $rst$ is reduced. If $rst\notin W^{I}$, then the nonreduced word $rsts=rt$ or $st$. The former is impossible since otherwise, we will again obtain that $m_{st}=2$, which is a contradiction. Hence $rsts=st$. It follows that the simple reflection $r\in W_{\{s,t\}}$, so $r=t$, and $m_{st}=3$. Using the correspondence between $RW(\sigma_{w})$ and $RW(\sigma_{w^{\prime}})$, the above argument can be applied to show that $m_{\tau(s)\tau(t)}=3$, so $a_{ts}=a^{\prime}_{\tau(t)\tau(s)}$ since $a_{st}=a^{\prime}_{\tau(s)\tau(t)}$ by the argument in the previous paragraph. Finally, if $rst\in W^{I}$, then Lemma 4.7 implies that
$\delta_{rt}-a_{rt}+a_{rs}a_{st}=\delta_{\tau(r)\tau(t)}-a^{\prime}_{\tau(r)\tau(t)}+a^{\prime}_{\tau(r)\tau(s)}a^{\prime}_{\tau(s)\tau(t)}.$
Since $\tau$ is bijective, $\delta_{rt}=\delta_{\tau(r)\tau(t)}$. Also, applying the argument in the previous paragraph again, we have $a_{rt}=a^{\prime}_{\tau(r)\tau(t)}$ and $a_{st}=a^{\prime}_{\tau(s)\tau(t)}\neq 0$. Hence $a_{rs}=a^{\prime}_{\tau(r)\tau(s)}$. In conclusion, we always have $a_{rs}=a^{\prime}_{\tau(r)\tau(s)}$. Note that the ordered pair $(t_{1},t_{2})=(r,s)$ or $(s,r)$, and the latter one can only appear when $rs=sr$. If $(t_{1},t_{2})=(r,s)$, then $(\tau(t_{1}),\tau(t_{2}))=(\tau(r),\tau(s))$, and hence $a_{t_{1}t_{2}}=a^{\prime}_{\tau(t_{1})\tau(t_{2})}$. Otherwise, $(t_{1},t_{2})=(s,r)$ and $(\tau(t_{1}),\tau(t_{2}))=(\tau(s),\tau(r))$. Since $sr\notin[1,w]^{I}$, $\tau(s)\tau(r)$ does not lie in $[1,w^{\prime}]^{I^{\prime}}$. Hence $\tau(s)\tau(r)=\tau(r)\tau(s)$. Therefore, $a_{t_{1}t_{2}}=-1=a^{\prime}_{\tau(t_{1})\tau(t_{2})}$. ∎
5. Constructing Isomorphism: Equal Supports
This section will study the sufficient conditions and prove the main theorem.
Let $A=(a_{st})_{(s,t)\in S^{2}}$ and $A^{\prime}=(a^{\prime}_{s^{\prime}t^{\prime}})_{(s^{\prime},t^{\prime})\in S^{\prime 2}}$ be two Cartan matrices with associated Weyl groups $W$ and $W^{\prime}$, and sets of simple reflections $S$ and $S^{\prime}$, respectively. Let $I\subset S$ and $I^{\prime}\subset S^{\prime}$. Take two fully supported elements $w\in W^{I}$ and $w^{\prime}\in W^{\prime I^{\prime}}$. We assume that there is a bijection $\tau:S\to S^{\prime}$ sending $I$ to $I^{\prime}$ such that:
  1. (a)
    for some reduced word $w=s_{1}\cdots s_{k}$, $w^{\prime}=\tau(s_{1})\cdots\tau(s_{k})$;
  2. (b)
    for any $t_{1},t_{2}\in S$, $a_{t_{1}t_{2}}=a^{\prime}_{\tau(t_{1})\tau(t_{2})}$ whenever $t_{1}t_{2}\leq w$.
We want to show that the varieties $X(w,A,\{s\})$ and $X(w^{\prime},A^{\prime},\{s^{\prime}\})$ are algebraically isomorphic. First, we have the following result.
Lemma 5.1.
The bijection $\tau:S\to S^{\prime}$ induces a poset isomorphism $[1,w]^{I}\to[1,w^{\prime}]^{I^{\prime}}$.
Proof.
If $v=s_{1}\cdots s_{k}\leq w$ is a reduced word, then we define $\tau(v)=\tau(s_{1})\cdots\tau(s_{k})$. By [RS21, Lemma 2.2], for any $v\in[1,w]$, $\tau(v)$ is well-defined, and $\tau$ induces a bijection
$RW(v)\to RW(\tau(v)),$ $(s_{1},\cdots,s_{k})\mapsto(\tau(s_{1}),\cdots,\tau(s_{k})),$
for any $v\in[1,w]$. It follows that $\tau$ also induces a bijection
$D_{R}(v)\to D_{R}(\tau(v)).$
[RS21, Lemma 2.4] implies that $\tau$ induces a poset isomorphism $[1,w]\to[1,w^{\prime}]$. Given $v\in[1,w]$, since
$v\in W^{I}\iff D_{R}(v)\cap I=\varnothing\iff D_{R}(\tau(v))\cap I^{\prime}=\varnothing\iff\tau(v)\in W^{I},$
the poset isomorphism $[1,w]\to[1,w^{\prime}]$ restricts to the isomorphism $[1,w]^{I}\to[1,w^{\prime}]^{I^{\prime}}$. ∎
In the following lemmas, we add the additional assumption that $a_{st}\leq a^{\prime}_{\tau(s)\tau(t)}$ for all simple reflections $s,t\in S$.
The bijection $\tau$ induces a bijection $\tau:\Delta\mapsto\Delta^{\prime}$ between simple roots, given by $\alpha_{s}\mapsto\alpha_{\tau(s)}$. Let $\mathfrak{g}=\mathfrak{n}^{-}(A)\oplus\mathfrak{h}(A)\oplus\mathfrak{n}^{+}(A)$ and $\mathfrak{g}(A^{\prime})=\mathfrak{n}^{-}(A^{\prime})\oplus\mathfrak{h}(A^{\prime})\oplus\mathfrak{n}^{+}(A^{\prime})$ be two complex reductive Lie algebras corresponding to $A$ and $A^{\prime}$, respectively. The lemma below is [RS21, Lemma 3.5 (a)].
Lemma 5.2.
There are surjective Lie algebra homomorphisms
$\varphi^{\pm}:\mathfrak{n}^{\pm}(A)\to\mathfrak{n}^{\pm}(A^{\prime}),$
given by $e_{\alpha}\mapsto e^{\prime}_{\tau(\alpha)}$ and $f_{\alpha}\mapsto f^{\prime}_{\tau(\alpha)}$, respectively.
The following lemma restates [RS21, Lemma 3.8].
Lemma 5.3.
The homomorphism $\varphi^{+}$ induces an isomorphism
$\varphi^{+}:\mathfrak{n}^{+}(A)_{v}\to\mathfrak{n}^{+}(A^{\prime})_{\tau(v)}$
for each $v\in[1,w]^{I}$.
If $\lambda$ and $\lambda^{\prime}$ are $I$ and $I^{\prime}$-regular weights, then we write $V=L_{\lambda}$ and $V^{\prime}=L_{\lambda^{\prime}}$ with corresponding highest weight vectors $\omega$ and $\omega^{\prime}$, respectively.
Lemma 5.4.
Let $\lambda$ and $\lambda^{\prime}$ be $I$ and $I^{\prime}$-regular weights, respectively. If $\lambda(h_{\alpha})=\lambda^{\prime}(h^{\prime}_{\tau(\alpha)})$ for each $\alpha\in\Delta$, then there is a surjective $\mathfrak{n}^{-}(A)$-homomorphism $\pi:V\to V^{\prime}$, sending the highest weight vector $\omega$ to $\omega^{\prime}$, where $V^{\prime}$ is regarded as a $\mathfrak{n}^{-}(A)$-module via the homomorphism $\varphi^{-}$. The homomorphism $\pi$ satisfies
$\pi(\exp(e)v\omega)=\exp(\varphi^{+}(e))\tau(v)\omega^{\prime}$
for all $e\in\mathfrak{n}^{+}(A)$ and $v\leq w$.
Remark.
The formula makes sense since the Weyl group element $v$ can be lifted to an element in the normalizer of the maximal torus, and different liftings acting on the vector $\omega$ give the same result.
Proof of Lemma 5.4.
The first part of the lemma is [RS21, Lemma 3.5 (b)]. The second part is [RS21, Lemma 3.9]. ∎
With the preparation above, we have the following sufficient conditions.
Theorem 5.5.
Let
$A=(a_{st})_{(s,t)\in S^{2}}$ and $A^{\prime}=(a^{\prime}_{s^{\prime}t^{\prime}})_{(s^{\prime},t^{\prime})\in S^{\prime 2}}$
be two Cartan matrices with associated Weyl groups $W$ and $W^{\prime}$, and sets of simple reflections $S$ and $S^{\prime}$, respectively. Let $I\subset S$ and $I^{\prime}\subset S^{\prime}$. Take $w\in W^{I}$ and $w^{\prime}\in W^{\prime I^{\prime}}$. If there is a bijection $\tau:S(w)\to S(w^{\prime})$ sending $S(w)\cap I$ to $S(w^{\prime})\cap I^{\prime}$ such that:
  1. (a)
    for some reduced word $w=s_{1}\cdots s_{k}$, $w^{\prime}=\tau(s_{1})\cdots\tau(s_{k})$ is also a reduced word;
  2. (b)
    for any $t_{1},t_{2}\in S$, $a_{t_{1}t_{2}}=a^{\prime}_{\tau(t_{1})\tau(t_{2})}$ whenever $t_{1}t_{2}\leq w$,
then $X(w,A,I)$ and $X(w^{\prime},A^{\prime},I^{\prime})$ are isomorphic.
Proof.
Using Lemma 3.9, we can suppose that $S(w)=S$ and $S(w^{\prime})=S^{\prime}$. First, we assume that $a_{st}\leq a^{\prime}_{s^{\prime}t^{\prime}}$. Take $\lambda$ and $\lambda^{\prime}$ to be $I$ and $I^{\prime}$-regular weights, respectively. By possibly increasing $\lambda$ or $\lambda^{\prime}$, we can assume that $\lambda(h_{\alpha})=\lambda^{\prime}(h^{\prime}_{\tau(\alpha)})$ for each $\alpha\in\Delta$. Take $G$ (resp. $P$) to be the reductive group (resp. the parabolic subgroup) corresponding to the flag variety $X(w,A)$. Then a Schubert cell can be uniquely written as $U_{v}vP/P$ for some $v\in W^{I}$, where $U_{v}$ is the unipotent subgroup with Lie algebra $\mathfrak{n}^{+}(A)_{v}$. Hence under the closed immersion given in Lemma 3.5, every element in $X(w,A,I)$ can be written uniquely as $[\exp(e)v\omega]\in\mathbb{P}(V)$, the line through $\exp(e)v\omega\in V$, for some $v\in[1,w]^{I}$ and $e\in\mathfrak{n}^{+}_{v}$. Then Lemma 5.3 and 5.4 show that $\pi$ restricts to a bijection between Schubert cells indexed by $v$ and $\tau(v)$. Hence Lemma 5.1 implies that $\pi$ restricts to a bijection
$X(w,A,I)\to X(w^{\prime},A^{\prime},I^{\prime}).$
Since Schubert varieties are normal, $X(w,A,I)$ and $X(w^{\prime},A^{\prime},I^{\prime})$ are isomorphic.
Next we remove the assumption $a_{st}\leq a^{\prime}_{\tau(s)\tau(t)}$. Then we define a Cartan matrix $A^{\prime\prime}=(a^{\prime\prime}_{st})$, where $a^{\prime\prime}_{st}=\max\{a_{st},a^{\prime}_{\tau(s)\tau(t)}\}$. Compared to the Dynkin diagram of $A$, the one of $A^{\prime\prime}$ is just the diagram obtained by removing several edges, so it is still of finite type. Based on the argument above, we conclude that
$X(w,A,I)\cong X(w,A^{\prime\prime},I)\cong X(w^{\prime},A^{\prime},I^{\prime}).$
Proof of Theorem 1.1.
Combining Lemma 3.9, Proposition 4.9, and Theorem 5.5, we conclude that conditions 1 and 2 in our main theorem are equivalent. The last statement in the main theorem follows from Proposition 4.5. ∎
As one may notice that the sufficient condition given in theorem 5.5 is a generalization of the one in the main theorem, one may conjecture that the necessary condition stated in Proposition 4.9 can also be generalized to all equally supported pairs.
Conjecture 5.6.
Let
$A=(a_{st})_{(s,t)\in S^{2}}$ and $A^{\prime}=(a^{\prime}_{s^{\prime}t^{\prime}})_{(s^{\prime},t^{\prime})\in S^{\prime 2}}$
be two Cartan matrices with associated Weyl groups $W$ and $W^{\prime}$, and sets of simple reflections $S$ and $S^{\prime}$, respectively. Let $I\subset S$ and $I^{\prime}\subset S^{\prime}$. Take $w\in W^{I}$ and $w^{\prime}\in W^{\prime I^{\prime}}$. Assume that $|S(w)|=|S(w^{\prime})|$. Then the following are equivalent:
  1. (1)
    the Schubert varieties $X(w,A,I)$ and $X(w^{\prime},A^{\prime},I^{\prime})$ are algebraically isomorphic;
  2. (2)
    there exists a bijection $\tau:S(w)\to S(w^{\prime})$ sending $S(w)\cap I$ to $S(w^{\prime})\cap I^{\prime}$, such that:
    1. (a)
      for some reduced word $w=s_{1}\cdots s_{k}$, $w^{\prime}=\tau(s_{1})\cdots\tau(s_{k})$ is also a reduced word;
    2. (b)
      for any $t_{1},t_{2}\in S(w)$, $a_{t_{1},t_{2}}=a^{\prime}_{\tau(t_{1}),\tau(t_{2})}$ whenever $t_{1}t_{2}\leq w$.
The main theorem can be deduced from Conjecture 5.6 as follows. Suppose that there exists a bijection $\sigma:S(w)\to S(w^{\prime})$ sending $S(w)\cap I$ to $S(w^{\prime})\cap I$ such that the conditions (a) and (b) in Theorem 1.1 hold. Then the assumption $|S(w)|=|S(w^{\prime})|$ in Conjecture 5.6 holds, so $X(w,A,I)$ and $X(w^{\prime},A^{\prime},I^{\prime})$ are isomorphic by the equivalence of conditions 1 and 2 given in Conjecture 5.6. On the other hand, suppose that $X(w,A,I)$ and $X(w^{\prime},A^{\prime},I^{\prime})$ are isomorphic. By Lemma 3.7, we have
$|S(w)\cap I|=|\Sigma_{1}(w,A,I)|=|\Sigma_{1}(w^{\prime},A^{\prime},I^{\prime})|=|S(w^{\prime})\cap I^{\prime}|.$
The identity $|S(w)\setminus I|=|S(w^{\prime})\setminus I^{\prime}|$, given by Proposition 4.5, implies that $|S(w)|=|S(w^{\prime})|$, so we may apply Conjecture 5.6 to finish the proof of Theorem 1.1.
Conjecture 5.6 holds in low-dimensional cases.
Proposition 5.7.
Conjecture 5.6 holds when the Schubert variety $X(w,A,I)$ is of dimension less than three.
Proof.
Since zero and one-dimensional Schubert varieties are isomorphic to a point and the projective line $\mathbb{P}^{1}$ respectively, the Conjecture 5.6 holds automatically. Now assume that $X(w,A,I)$ is of dimension two. Then $|S(w)|\leq\ell(w)=2$, so $|S(w)\cap I|\leq 1$. The statement follows from the main theorem. ∎
6. Folding by Automorphisms
There are many isomorphic pairs of Schubert varieties which are not equally supported. For instance, [RS21, Example 1.5] claims that the Schubert varieties
$X\left(s_{3}s_{2}s_{1},\begin{pmatrix}2&-1&0\\ -1&2&-1\\ 0&-1&2\end{pmatrix},\{s_{2},s_{3}\}\right)$ and $X\left(s_{1}s_{2}s_{1},\begin{pmatrix}2&-2\\ -1&2\end{pmatrix},\{s_{2}\}\right),$
of types $A_{3}$ and $C_{2}$ respectively, are algebraically isomorphic to the projective space $\mathbb{P}^{3}$. Thus we need another method to deal with such pairs. We consider the pairs obtained from folding the root system.
We recall the basic definitions and results of folding a root system.
[…]
$\langle\alpha_{\tau(s)},\alpha_{\tau(t)}\rangle=\langle\tau(\alpha_{s}),\tau(\alpha_{t})\rangle$
for both $s,t\in S$. If we extend the map $\alpha_{s}\mapsto\alpha_{\tau(s)}$ linearly, we may view $\tau$ as an isometry of $V$. Since $s_{\alpha}\tau s_{\alpha}^{-1}=s_{\tau(\alpha)}$, $\tau$ acts via conjugation as an automorphism of the Weyl group $W$. Denote ${\Delta_{i}}$ as the collection of $\tau$-orbits. The following assumes that the simple roots in the same $\tau$-orbit $\Delta_{i}$ are pairwise orthogonal. Let
$\beta_{i}^{\vee}=\sum_{\alpha\in\Delta_{i}}\alpha^{\vee}$
be the sum of simple coroots in the same $\tau$-orbit, and let $\Delta^{\tau}$ be the collection of all such $\beta_{i}$. Then $\Delta^{\tau}$ is a set of simple roots of some root system $R^{\tau}$. The root system $(R^{\tau},\Delta^{\tau})$ is called the root system obtained from folding the root system $(R,\Delta)$ by the automorphism $\tau$.
Figure 1 shows some examples of folding.
Figure 1. Examples of Folding
Denote
$t_{i}=\prod_{\alpha\in\Delta_{i}}s_{\alpha}\in W.$
Next two results are [St08, Claim 1] and [St08, Claim 3], respectively.
Lemma 6.1.
A sum of pairwise orthogonal coroots in a single $\tau$-orbit is a coroot of $R^{\tau}$. Conversely, all coroots of $R^{\tau}$ have this form.
Lemma 6.2.
The homomorphism $s_{\beta_{i}}\mapsto t_{i}$ is an isomorphism from the Weyl group of the root system $R^{\tau}$ to the subgroup $W^{\tau}$ of $W$ fixed by the action of $\tau$.
In other words, denoting the collection of $t_{i}=\prod_{\alpha\in\Delta_{i}}s_{\alpha}\in W$ by $S^{\tau}$, then $(W^{\tau},S^{\tau})$ is the Weyl of the root system $(R^{\tau},\Delta^{\tau})$.
Lemma 6.3.
Let
$A=(a_{s_{i}s_{j}})_{(s_{i},s_{j})\in S^{2}}$ and $A^{\tau}=(a_{t_{k}t_{l}})_{(t_{k},t_{l})\in{S^{\tau}}^{2}}$
be the Cartan matrices of the root systems $(R,\Delta)$ and $(R^{\tau},\Delta^{\tau})$, respectively. Then the formula
$a_{t_{i}t_{j}}=\sum_{s_{i}\in I_{i}}a_{s_{i}s_{j}}$
holds for any $s_{j}\in I_{j}$.
Proof.
By definition, we have
$\beta_{i}^{\vee}=\sum_{\alpha_{i}\in\Delta_{i}}\alpha_{i}^{\vee}$ and $\beta_{j}^{\vee}=\sum_{\alpha_{j}\in\Delta_{j}}\alpha_{j}^{\vee}.$
Recall that the simple roots in the same $\tau$-orbit $\Delta_{i}$ are pairwise orthogonal. We compute that
$a_{t_{i}t_{j}}$ $=\langle\beta_{i}^{\vee},\beta_{j}\rangle=\langle\beta_{i}^{\vee},\beta_{j}^{\vee\vee}\rangle=\frac{2\langle\beta_{i}^{\vee},\beta_{j}^{\vee}\rangle}{\langle\beta_{j}^{\vee},\beta_{j}^{\vee}\rangle}=\frac{2\left\langle\sum_{\alpha_{i}\in\Delta_{i}}\alpha_{i}^{\vee},\sum_{\alpha_{j}\in\Delta_{j}}\alpha_{j}^{\vee}\right\rangle}{\left\langle\sum_{\alpha_{j}\in\Delta_{j}}\alpha_{j}^{\vee},\sum_{\alpha_{j}\in\Delta_{j}}\alpha_{j}^{\vee}\right\rangle}$
$=\frac{2\sum_{\alpha_{i}\in\Delta_{i}}\sum_{\alpha_{j}\in\Delta_{j}}\langle\alpha_{i}^{\vee},\alpha_{j}^{\vee}\rangle}{\sum_{\alpha_{j}\in\Delta_{j}}\langle\alpha_{j}^{\vee},\alpha_{j}^{\vee}\rangle}=\frac{2\sum_{\alpha_{i}\in\Delta_{i}}\sum_{\alpha_{j}\in\Delta_{j}}\frac{2\langle\alpha_{i}^{\vee},\alpha_{j}\rangle}{\langle\alpha_{j},\alpha_{j}\rangle}}{\sum_{\alpha_{j}\in\Delta_{j}}\left\langle\frac{2\alpha_{j}}{\langle\alpha_{j},\alpha_{j}\rangle},\frac{2\alpha_{j}}{\langle\alpha_{j},\alpha_{j}\rangle}\right\rangle}$
$=\frac{\sum_{\alpha_{j}\in\Delta_{j}}\frac{1}{\langle\alpha_{j},\alpha_{j}\rangle}\sum_{\alpha_{i}\in\Delta_{i}}\langle\alpha_{i}^{\vee},\alpha_{j}\rangle}{\sum_{\alpha_{j}\in\Delta_{j}}\frac{1}{\langle\alpha_{j},\alpha_{j}\rangle}}.$
To complete the proof, it is sufficient to show that $\sum_{\alpha_{i}\in\Delta_{i}}\langle\alpha_{i}^{\vee},\alpha_{j}\rangle$ is independent on $\alpha_{i}\in\Delta_{i}$. In fact,
$\sum_{\alpha_{i}\in\Delta_{i}}\langle\alpha_{i}^{\vee},\alpha_{j}\rangle=\sum_{\alpha_{i}\in\Delta_{i}}\langle\tau(\alpha_{i}^{\vee}),\tau(\alpha_{j})\rangle=\sum_{\alpha_{i}\in\Delta_{i}}\langle\alpha_{i}^{\vee},\tau(\alpha_{j})\rangle.$
Lemma 6.4.
There is an injective Lie algebra homomorphism
$\varphi:\mathfrak{g}^{\tau}\to\mathfrak{g},$
given by $e_{t_{i}}\mapsto\sum_{s\in I_{i}}e_{s}$, $h_{t_{i}}\mapsto\sum_{s\in I_{i}}h_{s}$ and $f_{t_{i}}\mapsto\sum_{s\in I_{i}}f_{s}$, respectively.
Proof.
First we show that $\varphi$ is a Lie algebra homomorphism. It suffices to check the Serre’s relations; that is,
$\varphi([h_{t_{i}},h_{t_{j}}])$ $=0,$
$\varphi([h_{t_{i}},e_{t_{j}}])$ $=a_{t_{i}t_{j}}\varphi(e_{t_{j}}),$
$\varphi([h_{t_{i}},f_{t_{j}}])$ $=-a_{t_{i}t_{j}}\varphi(f_{t_{j}}),$
$\varphi([e_{t_{i}},f_{t_{j}}])$ $=\delta_{ij}h_{t_{i}},$
$\varphi\left((\mbox{ad }e_{t_{i}})^{1-a_{t_{i}t_{j}}}e_{t_{j}}\right)$ $=0,$
$\varphi\left((\mbox{ad }f_{t_{i}})^{1-a_{t_{i}t_{j}}}f_{t_{j}}\right)$ $=0.$
We compute them one by one. For the first one,
$\varphi([h_{t_{i}},h_{t_{j}}])=\left[\sum_{s_{i}\in I_{i}}h_{s_{i}},\sum_{s_{j}\in I_{j}}h_{s_{j}}\right]=\sum_{s_{i}\in I_{i}}\sum_{s_{j}\in I_{j}}[h_{s_{i}},h_{s_{j}}]=0.$
The second and the third ones are similar. In fact, we have
$\varphi([h_{t_{i}},e_{t_{j}}])=\sum_{s_{i}\in I_{i}}\sum_{s_{j}\in I_{j}}[h_{s_{i}},e_{s_{j}}]=\sum_{s_{i}\in I_{i}}\sum_{s_{j}\in I_{j}}a_{s_{i}s_{j}}e_{s_{j}}=\sum_{s_{j}\in I_{j}}a_{t_{i}t_{j}}e_{s_{j}}=a_{t_{i}t_{j}}\varphi(e_{t_{j}})$
and
$\varphi([h_{t_{i}},f_{t_{j}}])=-\sum_{s_{i}\in I_{i}}\sum_{s_{j}\in I_{j}}a_{s_{i}s_{j}}f_{s_{j}}=-a_{t_{i}t_{j}}\varphi(f_{t_{j}}).$
The fourth identity follows from the calculation
$\varphi([e_{t_{i}},f_{t_{j}}])=\sum_{s_{i}\in I_{i}}\sum_{s_{j}\in I_{j}}[e_{s_{i}},f_{s_{j}}]=\sum_{s_{i}\in I_{i}}\sum_{s_{j}\in I_{j}}\delta_{s_{i}s_{j}}h_{s_{i}}=\delta_{ij}\sum_{s_{i}\in I_{i}}h_{s_{i}}=\delta_{ij}\varphi(h_{t_{i}}).$
Now we show the fifth equation. Since we assume that the simple roots in the same $\tau$-orbit $\Delta_{i}$ are pairwise orthogonal, the operators ad $\mbox{ad }e_{s_{i}}$ commute with each other. Hence the left-hand side
$\varphi\left((\mbox{ad }e_{t_{i}})^{1-a_{t_{i}t_{j}}}e_{t_{j}}\right)=\sum_{s_{j}\in I_{j}}\left(\mbox{ad }\sum_{s_{i}\in I_{i}}e_{s_{i}}\right)^{1-a_{t_{i}t_{j}}}e_{s_{j}}$
is equal to zero by the pigeonhole principle and Serre’s relation
$(\mbox{ad }e_{s_{i}})^{1-a_{s_{i}s_{j}}}e_{s_{j}}=0.$
Similarly, the sixth equation also holds. Thus $\varphi$ is indeed a Lie algebra homomorphism.
By definition, $\varphi|_{\mathfrak{h}^{\tau}}:\mathfrak{h}^{\tau}\to\mathfrak{h}$ is injective. It remains to show that $\varphi|_{{\mathfrak{n}^{\tau}}^{\pm}}:{\mathfrak{n}^{\tau}}^{\pm}\to\mathfrak{n}^{\pm}$ is also injective. For any $\beta\in{R^{\tau}}^{+}$, it can be written as $\beta=t(\beta_{i})$ for some $t\in W^{\tau}$ and $\beta_{i}\in S^{\tau}$. If we write $\beta_{i}^{\vee}=\sum_{\alpha\in\Delta_{i}}\alpha^{\vee}$, then $\beta^{\vee}=\sum_{\alpha\in\Delta_{i}}t(\alpha^{\vee})$. Each coroot $t(\alpha^{\vee})\in R$ is positive since they lie in the same $\tau$-orbit, and $\tau$ sends positive (resp. negative) coroots to positive coroots (resp. negative). Again, by definition, the root space
$\mathfrak{g}^{\tau}_{\beta}=t(\mathfrak{g}^{\tau}_{\beta_{i}})\mapsto\sum_{\alpha\in\Delta_{i}}t(\mathfrak{g}_{\alpha})=\sum_{\alpha\in\Delta_{i}}\mathfrak{g}_{t(\alpha)},$
so the image of $\mathfrak{g}^{\tau}_{\beta}$ is a direct sum of root spaces corresponding to a $\tau$-orbit. Clearly, different $\beta\in{R^{\tau}}^{+}$ correspond to different orbits, so $\varphi|_{{\mathfrak{n}^{\tau}}^{+}}:{\mathfrak{n}^{\tau}}^{\pm}\to\mathfrak{n}^{+}$ is injective. Similarly, $\varphi|_{{\mathfrak{n}^{\tau}}^{-}}:{\mathfrak{n}^{\tau}}^{-}\to\mathfrak{n}^{-}$ is also injective. ∎
From the proof of Lemma 6.4, we conclude that every coroot of $R^{\tau}$ can be uniquely written as a sum of pairwise orthogonal coroots of $R$ in a single $\tau$-orbit. If $\beta^{\vee}=\sum_{\alpha}\alpha^{\vee}$, then we write the collection of such roots $\alpha$ as $\Delta_{\beta}$. It is well-defined.
7. Constructing Isomorphism: Folding
The theorem below provides a sufficient condition, enabling us to construct classes of isomorphic pairs from folding the root systems.
Theorem 7.1.
Let $R$ be a simply laced root system with simple roots $\Delta$, and let $\tau$ be a diagram automorphism of $(R,\Delta)$. The root system obtained from folding $(R,\Delta)$ by $\tau$ is denoted by $(R^{\tau},\Delta^{\tau})$. Assume the simple roots in the same $\tau$-orbit $\Delta_{i}$ are pairwise orthogonal. Let
$A=(a_{s_{i}s_{j}})_{(s_{i},s_{j})\in S^{2}}$ and $A^{\tau}=(a_{t_{k}t_{l}})_{(t_{k},t_{l})\in{S^{\tau}}^{2}}$
be the Cartan matrices of the root systems $(R,\Delta)$ and $(R^{\tau},\Delta^{\tau})$, respectively. Let $(W,S)$ be the Weyl group of $(R,\Delta)$, and identify the subgroup $(W^{\tau},S^{\tau})$ as the Weyl group of $(R^{\tau},\Delta^{\tau})$ as Lemma 6.2. Denote the $\tau$-orbit of $S$ to be $I_{i}$. Let $I\subset S$ and $I^{\tau}\subset S^{\tau}$ and take two elements $w\in W^{I}$ and $w^{\tau}\in{W^{\tau}}^{I^{\tau}}$. Assume that
  1. (a)
    for every $\tau$-orbit of $R$, there is at most one root does not lie in $R_{I}$;
  2. (b)
    $I^{\tau}=\{\prod_{s\in I_{i}}s\in W^{\tau}|I_{i}\subset I\}$;
  3. (c)
    the map
    $[1,w^{\tau}]^{I^{\tau}}\to[1,w]^{I},$ $v\mapsto v^{\min}$
    is bijective, where $v^{\min}$ is the minimal representative of $v$ in $W^{I}$,
then $X(w,A,I)$ and $X(w^{\tau},A^{\tau},I^{\tau})$ are isomorphic as algebraic varieties.
Suppose that every assumption in the statement of the theorem holds. We prove several lemmas.
Lemma 7.2.
The bijection $[1,w^{\tau}]^{I^{\tau}}\to[1,w]^{I}$ in condition (c) is a poset isomorphism preserving length; that is, for every reflection $v\in[1,w^{\tau}]^{I^{\tau}}$, the lengths $\ell_{W^{\tau}}(v)$ and $\ell_{W}(v^{\min})$ are the same.
Proof.
Lemma 3.2 implies that the bijection is order-preserving. To show it preserves length, we prove the following statement: for a reduced expression $t_{1}\cdots t_{k}\in[1,w^{\tau}]^{I^{\tau}}$, its image has the reduced expression $s_{1}\cdots s_{k}\in[1,w]^{I}$ with $s_{i}\in I_{i}$.
The case $k=0$ is trivial. Suppose that $k>0$ and $t_{1}\cdots t_{k}\in[1,w^{\tau}]^{I^{\tau}}$ are reduced expressions. Then $t_{2}\cdots t_{k}\in[1,w^{\tau}]^{I^{\tau}}$ is also a reduced expression. By induction, its image has the reduced expression $s_{2}\cdots s_{k}\in[1,w]^{I}$ with $s_{i}\in I_{i}$. Hence
$(t_{1}\cdots t_{k})^{\min}=\left(\prod_{s\in I_{1}}ss_{2}\cdots s_{k}\right)^{\min}>s_{2}\cdots s_{k}.$
It follows that $(t_{1}\cdots t_{k})^{\min}$ has the form $s_{1}\cdots s_{k}$ with $s_{1}\in I_{1}$, since it is the minimal element lying between $s_{2}\cdots s_{k}$ and $\prod_{s\in I_{1}}ss_{2}\cdots s_{k}$. This shows that our claim is true.
Using the claim, it is easy to see that the bijection is length-preserving. ∎
Let $\mathfrak{g}=\mathfrak{n}^{-}\oplus\mathfrak{h}\oplus\mathfrak{n}^{+}$ and $\mathfrak{g}^{\tau}={\mathfrak{n}^{\tau}}^{-}\oplus\mathfrak{h}^{\tau}\oplus{\mathfrak{n}^{\tau}}^{+}$ be two complex reductive Lie algebras corresponding to the root systems $(R,\Delta)$ and $(R^{\tau},\Delta^{\tau})$, respectively.
Lemma 7.3.
Let $v\in[1,w^{\tau}]^{I^{\tau}}$. Suppose that $\beta\in{R^{\tau}}^{+}$ and write $\beta^{\vee}=\sum_{\alpha\in\Delta_{\beta}}\alpha^{\vee}$. Then $\beta\in I_{W^{\tau}}(v)$ if and only if there exists a unique $\alpha\in\Delta_{\beta}$ such that $\alpha\in I_{W}(v^{\min})$.
Proof.
If $v^{-1}(\beta)\in{R^{\tau}}^{-}$, then $v^{-1}(\alpha)\in R^{-}$ for some $\alpha$. Since $v\in W^{\tau}$ is $\tau$-invariant, $v^{-1}(\alpha)\in R^{-}$ for every $\alpha$. If $v^{-1}(\alpha)\in R_{I}$ for all $\alpha$, then $v^{-1}(\beta)\in R^{\tau}\cap\mbox{Span }\Delta_{I}=R^{\tau}_{I^{\tau}}$ by condition (b). But $v\in{W^{\tau}}^{I^{\tau}}$ and $-v^{-1}(\beta)\in I_{W^{\tau}}(v^{-1})$, which contradicts Lemma 3.4. Hence there exists an $\alpha_{0}$ such that $v^{-1}(\alpha_{0})\in R^{-}\setminus R_{I}^{-}$, and it is unique by condition (a). In other words, $\alpha_{0}$ is the unique root lying in $I_{W}(v^{\min})$. ∎
By sending the root $\beta\in I_{W^{\tau}}(v)$ to the root $\alpha\in\Delta_{\beta}\cap I_{W}(v^{\min})$, we have an injection $I_{W^{\tau}}(v)\to I_{W}(v^{\min})$. By condition (d), it is bijective since the cardinality of the inversion set is just the length. We denote the image of $\beta\in I_{W^{\tau}}(v)$ by $\alpha_{\beta}$.
As we did in the equally supported case, if $\lambda$ and $\lambda^{\tau}$ are $I$ and $I^{\tau}$-regular weights, then we write $V=L_{\lambda}$ and $V^{\tau}=L_{\lambda^{\tau}}$ with corresponding highest weight vectors $\omega$ and $\omega^{\tau}$, respectively.
Lemma 7.4.
Let $\lambda$ and $\lambda^{\tau}$ be $I$ and $I^{\tau}$-regular weights, respectively. If $\lambda^{\tau}(h_{t_{i}})=\sum_{s\in I_{i}}\lambda(h_{s})$ for each $t_{i}\in S^{\tau}$, then there is a $\mathfrak{g}^{\tau}$-homomorphism $\pi:V^{\tau}\to V$, sending $\omega^{\tau}$ to $\omega$, where $V$ is regarded as an $\mathfrak{g}^{\tau}$-module via the homomorphism $\varphi$. The homomorphism $\pi$ satisfies
$\pi(\exp(e)v\omega^{\tau})=\exp(\varphi(e))v^{\min}\omega$
for all $e\in{\mathfrak{n}^{\tau}}^{+}$ and $v\in[1,w^{\tau}]^{I^{\tau}}$.
Proof.
Sending $\omega^{\tau}$ to $\omega$ induces an $\mathfrak{g}^{\tau}$-homomorphism $\pi_{0}:M_{\lambda^{\tau}}\to M_{\lambda}$. Recall that $V=M_{\lambda}/M_{\lambda}^{1}$ and $V^{\tau}=M_{\lambda^{\tau}}/M_{\lambda^{\tau}}^{1}$, where $M_{\lambda}^{1}$ and $M_{\lambda^{\tau}}^{1}$ are generated by $f_{s}^{\lambda(h_{s})+1}\omega$ and $f_{t}^{\lambda^{\tau}(h_{t})+1}\omega^{\tau}$, respectively. Notice that
$\pi_{0}\left(f_{t_{i}}^{\lambda^{\tau}(h_{t_{i}})+1}\omega^{\tau}\right)=\left(\sum_{s\in I_{i}}f_{s}\right)^{\sum_{s\in I_{i}}\lambda(h_{s})}\omega\in M_{\lambda}^{1}$
by pigeonhole principle. Hence $\pi_{0}$ induces an ${\mathfrak{n}^{\tau}}^{-}$-homomorphism $\pi:V^{\tau}\to V$, as required.
Now we prove the formula. First, both the action of $e\in{\mathfrak{n}^{\tau}}^{+}$ and the action $f\in{\mathfrak{n}^{\tau}}^{-}$ on $V^{\tau}$, so we have
$\pi(\exp(e)\eta_{1})=\exp(\varphi(e))\pi(\eta_{1})$ and $\pi(\exp(f)\eta_{1})=\exp(\varphi(f))\pi(\eta_{1})$
for any $\eta_{1}\in V^{\tau}$. The simple reflection $t_{i}\in S^{\tau}$ on $V$ acts the same as the action of $\exp(f_{t_{i}})\exp(-e_{t_{i}})\exp(f_{t_{i}})$. Applying the above formulas, for each $\eta_{2}\in V^{\tau}$, we compute that
$\pi(t_{i}\eta_{2})$ $=\pi(\exp(f_{t_{i}})\exp(-e_{t_{i}})\exp(f_{t_{i}})\eta_{2})=\exp(\varphi(f_{t_{i}}))\exp(\varphi(-e_{t_{i}}))\exp(\varphi(f_{t_{i}}))\pi(\eta_{2})$
$=\exp\left(\sum_{s_{i}\in I_{i}}f_{s_{i}}\right)\exp\left(-\sum_{s_{i}\in I_{i}}e_{s_{i}}\right)\exp\left(\sum_{s_{i}\in I_{i}}f_{s_{i}}\right)\pi(\eta_{2})$
$=\prod_{s_{i}\in I_{i}}\exp(f_{s_{i}})\prod_{s_{i}\in I_{i}}\exp(-e_{s_{i}})\prod_{s_{i}\in I_{i}}\exp(f_{s_{i}})\pi(\eta_{2})$
$=\prod_{s_{i}\in I_{i}}(\exp(f_{s_{i}})\exp(-e_{s_{i}})\exp(f_{s_{i}}))\pi(\eta_{2})=\prod_{s_{i}\in I_{i}}s_{i}\pi(\eta_{2})=t_{i}\pi(\eta_{2}).$
Thus, the $\mathfrak{g}^{\tau}$-homomorphism $\pi$ commutes with the Weyl group $W^{\tau}$-action. Combining these two results, we conclude that
$\pi(\exp(e)v\omega^{\tau})=\exp(\varphi(e))v\omega=\exp(\varphi(e))v^{\min}\omega$
since $\omega$ is the highest weight vector corresponding to the $I$-regular weight $\lambda$. ∎
For $v\in[1,w^{\tau}]^{I^{\tau}}$, we denote the root space decompositions
${\mathfrak{n}^{\tau}_{v}}^{+}=\bigoplus_{\beta\in I_{W^{\tau}}(v)}\mathfrak{g}^{\tau}_{\beta}$ and $\mathfrak{n}_{v^{\min}}^{+}=\bigoplus_{\alpha\in I_{W}(v^{\min})}\mathfrak{g}_{\alpha}.$
Lemma 7.5.
Let $v\in[1,w^{\tau}]^{I^{\tau}}$. The homomorphism $\pi$ defines a bijection
$\exp({\mathfrak{n}^{\tau}_{v}}^{+})v\omega^{\tau}\mapsto\exp(\mathfrak{n}_{v^{\min}}^{+})v^{\min}\omega.$
Proof.
First, we show that $\pi(\exp({\mathfrak{n}^{\tau}_{v}}^{+})v\omega^{\tau})\subset\exp(\mathfrak{n}_{v^{\min}}^{+})v^{\min}\omega$. Take a nonzero root vector $e_{\beta}\in\mathfrak{g}^{\tau}_{\beta}$. Then $\varphi(e_{\beta})=\sum_{\alpha\in\Delta_{\beta}}e_{\alpha}$ with $e_{\alpha}\in\mathfrak{g}_{\alpha}$ by the proof of Lemma 6.4. Then for constants $a_{\beta}$, we compute that
$\exp\left(\sum_{\beta\in I_{W^{\tau}}(v)}a_{\beta}e_{\beta}\right)v\omega^{\tau}\mapsto\exp\left(\sum_{\beta\in I_{W^{\tau}}(v)}a_{\beta}\sum_{\alpha\in\Delta_{\beta}}e_{\alpha}\right)v^{\min}\omega$
$=$ $\exp\left(\sum_{\alpha\in I_{W}(v^{\min})}a_{\alpha}e_{\alpha}+\sum_{\beta\in I_{W^{\tau}}(v)}a_{\beta}\sum_{\alpha\in\Delta_{\beta}\setminus\{\alpha_{\beta}\}}e_{\alpha}\right)v^{\min}\omega,$
by Lemma 7.3. Here, we write $a_{\alpha}=a_{\alpha_{\beta}}$ for simplicity. We denote the exponential above by $E$. Let $\alpha_{1}\in R^{+}\setminus I_{W}(v^{\min})$ be the root appearing in the exponential above with minimal height and let $c_{1}$ be its coefficient. Then any root in $R^{+}\setminus I_{W}(v^{\min})$ appearing in $E\exp(-c_{1}e_{\alpha_{1}})$ (writing it using a single exponential according to the Campbell-Hausdorff formula) has height no less than the height of $\alpha_{1}$, and the number of roots with the same height as $\alpha_{1}$ is strictly less, which can be seen by using the Campbell-Hausdorff formula. Applying the algorithm several times, the minimal height of the roots appearing in $E\exp(-c_{1}e_{\alpha_{1}})\cdots\exp(-c_{k}e_{\alpha_{k}})$ is strictly larger than the height of $\alpha_{1}$. Therefore, using the algorithm continuously, there will not be any root appearing in $E\exp(-c_{1}e_{\alpha_{1}})\cdots\exp(-c_{n}e_{\alpha_{n}})=:E^{\prime}$ lying in $R^{+}\setminus I_{W}(v^{\min})$, and each $\alpha_{i}\in R^{+}\setminus I_{W}(v^{\min})$, since the height of a root is bounded. Therefore,
$Ev^{\min}\omega$ $=E^{\prime}\exp(c_{n}e_{\alpha_{n}})\cdots\exp(c_{1}e_{\alpha_{1}})v^{\min}\omega$
$=v^{\min}\cdot{v^{\min}}^{-1}(E^{\prime})\exp\left(c_{n}e_{{v^{\min}}^{-1}(\alpha_{n})}\right)\cdots\exp\left(c_{1}e_{{v^{\min}}^{-1}(\alpha_{1})}\right)\omega$
$=v^{\min}\cdot{v^{\min}}^{-1}(E^{\prime})\omega=E^{\prime}v^{\min}\omega\in\exp(\mathfrak{n}_{v^{\min}}^{+})v^{\min}\omega,$
since $\omega$ is an $I$-regular weight vector.
Now we show that the restriction of $\pi$ is bijective. Under the algorithm, the coefficient of $e_{\alpha}$ for $\alpha\in I_{W}(v^{\min})$ becomes a polynomial $f_{\alpha}$. Hence there is a morphism between the affine spaces
$\mathfrak{n}_{v^{\min}}^{+}\to\mathfrak{n}_{v^{\min}}^{+},$ $(a_{\alpha})\mapsto(g_{\alpha}).$
The algorithm is invertible, and its inverse also turns the coefficient $b_{\alpha}$ of $e_{\alpha}$ ($\alpha\in I_{W}(v^{\min})$) into a polynomial $g_{\alpha}$. Thus, we have the inverse morphism
$\mathfrak{n}_{v^{\min}}^{+}\to\mathfrak{n}_{v^{\min}}^{+},$ $(b_{\alpha})\mapsto(h_{\alpha}).$
Therefore, $\pi$ restricts to a bijection
$\exp({\mathfrak{n}^{\tau}_{v}}^{+})v\omega^{\tau}\mapsto\exp(\mathfrak{n}_{v^{\min}}^{+})v^{\min}\omega,$ $\exp\left(\sum_{\beta\in I_{W^{\tau}}(v)}a_{\beta}e_{\beta}\right)v\omega^{\tau}\mapsto\exp\left(\sum_{\alpha\in I_{W}(v^{\min})}g_{\alpha}e_{\alpha}\right)v^{\min}\omega.$
Proof of Theorem 7.1.
Take $\lambda$ and $\lambda^{\tau}$ to be $I$ and $I^{\tau}$-regular weights, respectively, and assume that $\lambda(h_{\alpha})=\lambda^{\prime}(h^{\prime}_{\tau(\alpha)})$ for each $\alpha\in\Delta$. Take $G$ (resp. $P$) to be the reductive group (resp. the parabolic subgroup) corresponding to the flag variety $X(w^{\tau},A^{\tau},I^{\tau})$. Then a Schubert cell can be uniquely written as $U_{v}vP/P$ for some $v\in{W^{\tau}}^{I^{\tau}}$, where $U_{v}$ is the unipotent subgroup with Lie algebra ${\mathfrak{n}^{\tau}}^{+}_{v}$. By Lemma 3.5, every element in $X(w^{\tau},A^{\tau},I^{\tau})$ can be written uniquely as $[\exp(e)v\omega]\in\mathbb{P}(V)$, the line through $\exp(e)v\omega\in V$, for some $v\in[1,w^{\tau}]^{I^{\tau}}$ and $e\in{\mathfrak{n}^{\tau}}^{+}_{v}$. Then Lemma 7.4 yields that the $\mathfrak{g}^{\tau}$-homomorphism $\pi$ defines a morphism from $X(w^{\tau},A^{\tau},I^{\tau})$ to $X(w,A,I)$. By Lemma 7.5, $\pi$ is bijective on each pair of Borel cells $\exp({\mathfrak{n}^{\tau}_{v}}^{+})v\omega^{\tau}$ and $\exp(\mathfrak{n}_{v^{\min}}^{+})v^{\min}\omega$. By condition (c), this gives the bijection between Schubert varieties $X(w^{\tau},A^{\tau},I^{\tau})$ and $X(w,A,I)$. Therefore, the normality of Schubert varieties implies that $X(w^{\tau},A^{\tau},I^{\tau})$ are $X(w,A,I)$ isomorphic as algebraic varieties. ∎
8. Schubert Varieties in flag Varieties Corresponding to Maximal Parabolic Subgroups
Suppose that $(R,\Delta)$ is a simply laced root system with the Weyl group $(W,S)$ and $\tau$ is a diagram automorphism. The root system $(R^{\tau},\Delta^{\tau})$ with the Weyl group $(W^{\tau},S^{\tau})$ is obtained from folding by $\tau$. Choosing $I\subset S$ and $I^{\tau}\subset S^{\tau}$ as two proper subsets. Assume that they satisfy the conditions in Theorem 7.1. Without loss of generality, we can suppose that $(R^{\tau},\Delta^{\tau})$ is irreducible since any Schubert variety from a reducible root system is a product of Schubert varieties from irreducible root systems. Throughout this section, we use the type of a Cartan matrix to indicate the matrix itself unless there is ambiguity.
If $(R^{\tau},\Delta^{\tau})$ is still simply laced, then $(R,\Delta)$ is reducible, and each component can be identified as $(R^{\tau},\Delta^{\tau})$, so we may identify the Weyl group $W$ as ${W^{\tau}}^{n}$, where $n$ is the number of components. The action of automorphism $\tau$ is simply identifying each component as $(R^{\tau},\Delta^{\tau})$. Suppose that $\alpha\notin R_{I}$ is the highest root in one of the irreducible components of $(R,\Delta)$. Then any other root in the $\tau$-orbit containing $\alpha$ must lie in $R_{I}$. Hence any roots in other irreducible components of $(R,\Delta)$ lie in $R_{I}$. Then we have
$W^{I}={W^{\tau}}^{I^{\tau}}\times({W^{\tau}}^{S^{\tau}})^{n}={W^{\tau}}^{I^{\tau}}.$
Thus, the isomorphism given in Theorem 7.1 is trivial.
Example 8.1.
Consider the pair of root systems $A_{n}\times A_{n}$ and $A_{n}$ (see Figure 2). Write $S=\{s_{1},\cdots s_{n},s^{\prime}_{1},\cdots s^{\prime}_{n}\}$ and $S^{\tau}=\{t_{1},\cdots t_{n}\}$. We take $I=\{s^{\prime}_{1},\cdots s^{\prime}_{n}\}$, for example. If $w\in W^{I}$, then $s^{\prime}_{i}\notin S(w)$. Thus if we write $w=s_{i_{1}}\cdots s_{i_{k}}\in W^{I}$, then $w^{\tau}=t_{i_{1}}\cdots t_{i_{k}}\in{W^{\tau}}^{I^{\tau}}$. Let $G=SL_{n+1}(\mathbb{C})$, $B$ the standard Borel subgroup, and $P=B\times G$ the parabolic subgroup of type $A_{n}\times A_{n}$ corresponding to $I$. Then
$X(w,A,I)=(B\times B)wP/P\cong BwB/B\times BG/G\cong BwB/B=X(w^{\tau},A^{\tau},I^{\tau}).$
Figure 2. The Pair of Root Systems $A_{n}\times A_{n}$ and $A_{n}$
Next, we assume that $(R^{\tau},\Delta^{\tau})$ is not simply laced. Restricting $\tau$ to a component of $(R,\Delta)$ still gives an automorphism satisfying the conditions in Theorem 7.1, so we may suppose that $(R,\Delta)$ is irreducible. There are exactly four cases. We study them one by one.
Example 8.2.
Consider the pair of root systems $A_{2n-1}$ and $C_{n}$ (see Figure 3). The roots $\alpha_{2}+\cdots+\alpha_{2n-1}$ and $\alpha_{1}+\cdots+\alpha_{2n-2}$ lie on the same $\tau$-orbit, so two of them are contained in $R^{I}$. We can therefore assume that $s_{1}\in S$ (resp. $t_{1}\in S^{\tau}$) is the unique simple reflection not contained in $I$ (resp. $I^{\tau}$). Both the posets $W^{I}$ and ${W^{\tau}}^{I^{\tau}}$ are chains, whose poset structures are given by
$W^{I}=\{1<s_{1}<s_{2}s_{1}<\cdots<s_{2n-1}s_{2n-2}\cdots s_{1}\}$
and
${W^{\tau}}^{I^{\tau}}=\{1<t_{1}<t_{2}t_{1}<\cdots<t_{n}\cdots t_{1}<t_{n-1}t_{n}\cdots t_{1}<t_{1}\cdots t_{n}\cdots t_{1}\}.$
Hence the canonical map ${W^{\tau}}^{I^{\tau}}\to W^{I}$ is a bijection. By Theorem 7.1 and Lemma 3.9, we have the isomorphism
$X(t_{k}\cdots t_{1},C_{n},S^{\tau}\setminus\{t_{1}\})$ $\cong X(s_{k}\cdots s_{1},A_{2n-1},S\setminus\{s_{1}\}),$
$X(t_{k}\cdots t_{n}\cdots t_{1},C_{n},S^{\tau}\setminus\{t_{1}\})$ $\cong X(s_{2n-k}\cdots s_{n}\cdots s_{1},A_{2n-1},S\setminus\{s_{1}\}),$
for $k\leq n$.
Figure 3. The Pair of Root Systems $A_{2n-1}$ and $C_{n}$
Example 8.3.
Consider the pair of root systems $D_{n+1}$ and $B_{n}$ (see Figure 4). The roots $\alpha_{1}+\cdots+\alpha_{n}$ and $\alpha_{1}+\cdots+\alpha_{n-1}+\alpha_{n+1}$ lie on the same $\tau$-orbit, so we can assume that $s_{n+1}\in S$ (resp. $t_{n}\in S^{\tau}$) is the unique simple reflection not contained in $I$ (resp. $I^{\tau}$). The poset structures of $W^{I}$ and ${W^{\tau}}^{I^{\tau}}$ are messy, so we will not give the complete list of isomorphisms, but only give an example of the pair $D_{4}$ and $B_{3}$. The poset structures of $[1,t_{1}t_{3}t_{2}t_{3}]^{I^{\tau}}$ and $[1,s_{1}s_{3}s_{2}s_{4}]^{I}$ are given below,
${{t_{1}t_{3}t_{2}t_{3}}}$${{s_{1}s_{3}s_{2}s_{4}}}$${{t_{1}t_{2}t_{3}}}$${{t_{3}t_{2}t_{3}}}$${{s_{1}s_{2}s_{4}}}$${{s_{3}s_{2}s_{4}}}$${{t_{2}t_{3}}}$${{s_{2}s_{4}}}$${{t_{3}}}$${{s_{4}}}$11{1}111{1}1
which are identical. Hence Theorem 7.1 implies the isomorphism
$X(t_{1}t_{3}t_{2}t_{3},B_{3},S^{\tau}\setminus\{t_{4}\})\cong X(s_{1}s_{3}s_{2}s_{4},D_{4},S\setminus\{s_{4}\}).$
Figure 4. The Pair of Root Systems $D_{n+1}$ and $B_{n}$
Example 8.4.
Consider the pair of root systems $D_{4}$ and $G_{2}$ (see Figure 5). The roots $\alpha_{i}+\alpha_{4}$ ($i=1,2,3$) lie on the same $\tau$-orbit, so we assume that $s_{1}\in S$ is the unique simple reflection not contained in $I$, and hence $I^{\tau}=\{t_{2}\}$. In $G_{2}$, we have $t_{1}<t_{2}t_{1}<t_{1}t_{2}t_{1}$ in ${W^{\tau}}^{I^{\tau}}$. But
$t_{1}t_{2}t_{1}=(s_{1}s_{2}s_{3})s_{4}(s_{1}s_{2}s_{3})=(s_{2}s_{3})(s_{1}s_{4}s_{1})(s_{2}s_{3})=(s_{2}s_{3})(s_{4}s_{1}s_{4})(s_{2}s_{3})\mapsto s_{2}s_{3}s_{4}s_{1}\in W^{I},$
and $t_{1}t_{2}t_{1}\in{W^{\tau}}^{I^{\tau}}$, $s_{2}s_{3}s_{4}s_{1}\in W^{I}$ having different lengths. Hence Theorem 7.1 only gives two isomorphisms in this case, which are
$X(s_{1},A,I)\cong X(t_{1},A^{\tau},I^{\tau})$ and $X(s_{4}s_{1},A,I)\cong X(t_{2}t_{1},A^{\tau},I^{\tau}).$
The first isomorphism is trivial since any Schubert curve is isomorphic to the projective line $\mathbb{P}^{1}$. For the second one, the Schubert surfaces above are isomorphic to the projective plane $\mathbb{P}^{2}$ by the discussion in Example 2.2.
Figure 5. The Pair of Root Systems $D_{4}$ and $G_{2}$
Example 8.5.
Consider the pair of root systems $E_{6}$ and $F_{4}$ (see Figure 6). The roots $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}$ and $\alpha_{3}+\alpha_{4}+\alpha_{5}+\alpha_{6}$ lie on the same $\tau$-orbit, so we assume that $s_{1}\in S$ (resp. $t_{1}\in S^{\tau}$) is the unique simple reflection not contained in $I$ (resp. $I^{\tau}$). The canonical map ${W^{\tau}}^{I^{\tau}}\to W^{I}$ gives more isomorphisms by Theorem 7.1.
Figure 6. The Pair of Root Systems $E_{6}$ and $F_{4}$
Now we turn to studying the Schubert threefolds. A Schubert threefold is a Schubert variety of dimension three.
Theorem 8.6.
There are at most $34$ isomorphism classes of Schubert threefolds.
Proof.
By Lemma 3.9, it is sufficient to consider the Schubert threefolds $(w,A,I)$ with fully supported $w$. Then the number of simple reflections in $I$ is at most three.
First, we consider the Schubert threefolds from reducible root systems. Then it is obvious that such a Schubert threefold is a product of a Schubert surface and a projective line $\mathbb{P}^{1}$ since any Schubert curve is isomorphic to $\mathbb{P}^{1}$. By Theorem 2.6, there are 7 isomorphism classes.
Next, suppose that the Schubert threefolds $(w,A,I)$ come from irreducible root systems. If $A$ has rank 2, then it has type $A_{2}$, $B_{2}$, or $G_{2}$. As what we did in Section 2, for $n=1,2,3$, we denote
$M_{n}=\begin{pmatrix}2&-1\\ -n&2\end{pmatrix}.$
By Theorem 1.1, any two of the 10 Schubert threefolds
$X(s_{1}s_{2}s_{1},M_{1},\varnothing),$ $X(s_{1}s_{2}s_{1},M_{1},\{s_{2}\}),$
$X(s_{1}s_{2}s_{1},M_{2},\varnothing),X(s_{2}s_{1}s_{2},M_{2},\varnothing),$ $X(s_{1}s_{2}s_{1},M_{2},\{s_{2}\}),X(s_{1}s_{2}s_{1},M_{2},\{s_{1}\}),$
$X(s_{1}s_{2}s_{1},M_{3},\varnothing),X(s_{2}s_{1}s_{2},M_{3},\varnothing),$ $X(s_{1}s_{2}s_{1},M_{3},\{s_{2}\}),X(s_{1}s_{2}s_{1},M_{3},\{s_{1}\})$
are not isomorphic to each other.
Finally, we study the Schubert threefolds $(w,A,I)$ from irreducible root systems of rank 3. Then $A$ has type $A_{3}$, $B_{3}$, or $C_{3}$. We denote the Cartan matrices
$N_{1}=\begin{pmatrix}2&-1&0\\ -1&2&-1\\ 0&-1&2\end{pmatrix},$ $N_{2}=\begin{pmatrix}2&-1&0\\ -1&2&-1\\ 0&-2&2\end{pmatrix},$ $N_{3}=\begin{pmatrix}2&-1&0\\ -1&2&-2\\ 0&-1&2\end{pmatrix}.$
The element $w\in W^{I}$ can be $s_{1}s_{2}s_{3}$, $s_{1}s_{3}s_{2}$, $s_{2}s_{1}s_{3}$, or $s_{3}s_{2}s_{1}$. If $I$ is empty, then there are 6 isomorphism classes
$X(s_{1}s_{2}s_{3},N_{1},\varnothing),$ $X(s_{1}s_{2}s_{3},N_{3},\varnothing),$
$X(s_{1}s_{3}s_{2},N_{1},\varnothing),$ $X(s_{1}s_{3}s_{2},N_{2},\varnothing),$
$X(s_{2}s_{1}s_{3},N_{1},\varnothing),$ $X(s_{2}s_{1}s_{3},N_{3},\varnothing).$
Here, we note that there are isomorphisms
$X(s_{1}s_{2}s_{3},N_{1},\varnothing)\cong X(s_{3}s_{2}s_{1},N_{1},\varnothing)$ and $X(s_{1}s_{2}s_{3},N_{3},\varnothing)\cong X(s_{3}s_{2}s_{1},N_{2},\varnothing).$
If $I$ is a singleton, then there are 8 isomorphism classes
$X(s_{1}s_{2}s_{3},N_{1},\{s_{1}\}),X(s_{1}s_{2}s_{3},N_{3},\{s_{1}\}),$ $X(s_{1}s_{2}s_{3},N_{1},\{s_{2}\}),X(s_{1}s_{2}s_{3},N_{3},\{s_{2}\}),$
$X(s_{1}s_{3}s_{2},N_{1},\{s_{1}\}),$ $X(s_{1}s_{3}s_{2},N_{2},\{s_{1}\}),$
$X(s_{2}s_{1}s_{3},N_{1},\{s_{2}\}),$ $X(s_{2}s_{1}s_{3},N_{3},\{s_{2}\}).$
If $|I|=2$, then there are at most 4 isomorphism classes
$X(s_{1}s_{2}s_{3},N_{1},\{s_{1},s_{2}\}),$ $X(s_{1}s_{2}s_{3},N_{3},\{s_{1},s_{2}\}),$
$X(s_{1}s_{3}s_{2},N_{1},\{s_{1},s_{3}\}),$ $X(s_{1}s_{3}s_{2},N_{2},\{s_{1},s_{3}\}).$
From Example 8.2 or [RS21, Example 1.5], we know that the Schubert varieties
$X(s_{1}s_{2}s_{3},N_{1},\{s_{1},s_{2}\})$ and $X(s_{2}s_{1}s_{2},M_{2},\{s_{1}\})$
are isomorphic. Hence there are at most $7+10+6+8+4-1=34$ Schubert threefolds up to isomorphism. ∎
[…]